### v2

# Spirals and Beyond:

Abstract
We consider the problem of minimizing the worst-case search time for a hidden point target in the plane using multiple mobile agents of differing speeds, all starting from a common origin. The search time is normalized by the target’s distance to the origin, following the standard convention in competitive analysis. The objective is to minimize the maximum such normalized time over all possible target locations, known as the search cost. As a foundation for our main results, we extend the known result for a single unit-speed agent, which achieves a provably optimal search cost of approximately $\mathcal{U}_{1}=17.28935$ via a logarithmic spiral, to the setting of $n$ unit-speed agents. We design a symmetric spiral-based algorithm in which each agent follows a logarithmic spiral offset by equal angular phases. A key feature of this construction is that the resulting search cost is independent of which agent finds the target. We provide a closed-form upper bound $\mathcal{U}_{n}$ for the search cost in this setting, which forms the basis for our general result.
Our main contribution is an upper bound on the worst-case normalized search time for $n$ agents with arbitrary speeds. Inspired by natural optimality conditions, we describe an algorithmic framework that selects a subset of the agents and assigns spiral-type trajectories with speed-dependent angular offsets, ensuring that the search cost is independent of which agent reaches the target. A key corollary shows that a team of $n$ multi-speed agents, with the fastest agent having speed 1, can outperform $k$ unit-speed agents (i.e., achieve cost below $\mathcal{U}_{k}$) if the geometric mean of their speeds exceeds $\mathcal{U}_{n}/\mathcal{U}_{k}$. An implication of this result is that slow agents may be excluded from the spiral search if their presence lowers the geometric mean too much. This motivates the study of non-spiral algorithms that can do better in such cases. We establish new upper bounds for point search in cones and conic complements using a single unit-speed agent, which may be of independent interest. These bounds are then used to design hybrid strategies that combine spiral and directional search, and we show that these can outperform the previously derived spiral-based algorithms when some agents are sufficiently slow. This suggests that spiral trajectories may not be optimal in the general multi-speed setting.
1 Introduction
The lost-in-a-forest problem, introduced by Bellman in 1955 [16], asks how to navigate from an unknown location in a known region to its boundary while minimising expected travel time. It remains open and foundational to search theory. Many variants have since been studied, yielding notable upper bounds and generally loose lower bounds. We review selected results in Section 1.1.
A central variant involves a unit-speed agent searching for a hidden shoreline in the plane, modelled as an infinite line. The case where the shoreline distance is known was solved in 1957 [60], one of the few known optimal results. The unknown-distance case remains open, with only conjectured optimal upper bounds [8] and weak unconditional lower bounds [10]. The best strategy known uses a logarithmic spiral, and the best lower bound known [68] is not tight.
We consider a related variant introduced in [49], where the target is a point in the plane. A point is exposed by an agent if it lies on the line segment between the agent’s current position and the origin. The goal is to design a trajectory that minimises the worst-case target exposure time, normalized by the target’s distance from the origin. The logarithmic spiral was first shown to be optimal among monotone periodic trajectories [49], and later proven unconditionally optimal in [67].
We extend the upper bound results to a team of multi-speed agents, a rarely studied generalization. First, we construct natural, non-intersecting logarithmic spiral trajectories where all agents incur equal cost, and identify speed regimes where agents do or do not contribute to exposing new points. As a corollary, we relate multi-speed performance to that of unit-speed agents, and give a condition, based on the geometric mean of the fastest agents’ speeds, for effective contribution. Second, motivated by the limitations of spiral trajectories in low-speed regimes, we introduce and analyze non-spiral algorithms and show that they can outperform the previously derived spirals in such settings. This suggests that spiral-search may not always be optimal, as has been shown or conjectured in other models. Throughout the paper, statements of optimality are with respect to the underlying point search problem as defined above, the standard performance measure, and the agent speed regimes explicitly specified in each result.
1.1 Related Work
Search and exploration problems have a long history in operations research, robotics, and theoretical computer science. Early formal models date back to Koopman [63], with foundational work on deterministic and probabilistic search by Beck and Bellman [14, 18]. Stone [76] later introduced Bayesian models. This early work evolved into modern search theory, summarized in various surveys and books [5, 6, 25, 32, 49, 64]. A core concept nowadays in search theory is the competitive ratio [21], defined as the worst-case ratio between the cost of an online algorithm (without full information) and that of an optimal offline strategy. This and related frameworks [71] have led to extensive work on both deterministic [4, 3, 37] and randomized or fault-tolerant search [33, 39].
In one dimension, the classic Cow-Path problem asks a single agent to optimally search for a point. A spiral-like, exponentially expanding strategy is optimal [8]. Numerous variants have been studied, including cost-based [37], geometric [35], and agent-specific models [26, 51], all employing cyclic, monotonic spirals. An exception arises in the case of probabilistically faulty agents on a half-line, where monotone spirals are suboptimal [20].
In two dimensions, a closely related problem is shoreline search, introduced by Bellman [17], where the target is an infinite line. The case with known shoreline distance to the agent’s starting position was solved by Isbell [60], while the general “lost-in-a-forest” version (where the shape of the forest is part of the input) remains open [19, 45], with specific convex environments and general shapes having also been studied [56, 65]. For unknown distances, the best strategy known for shoreline search is a logarithmic spiral with competitive ratio $13.81$ [8, 46] and the best lower bounds known are $6.3972$ (unconditional) [10] and $12.5385$ (for cyclic trajectories) [68].
When the target is a point, which is the focus of this work, a logarithmic spiral achieves a search cost of $17.289$, and has been proven to be optimal [67]. This result resolved a long-standing conjecture that spiral trajectories are best among all possible strategies. Spiral strategies in planar domains have been extensively studied [47, 46]. Gal [49] established optimality within the class of monotone, periodic strategies, and Langetepe [67] later proved unconditional optimality, confirming the conjecture. For known-distance variants, recent work provides exact and average-case optimal results [27]. Related geometric search models also admit non-spiral optimal trajectories; for example, recent work based on Fermat’s principle establishes optimal average-case strategies for disk inspection in a different setting [55].
Collaborative search with multiple agents has also received significant attention, see [2, 23, 33] and [36, 38, 39, 59, 70]. The well-known ANTS problem [43] considers identical agents searching for an adversarially placed treasure in the plane, with performance depending on agent count and target distance. This extends earlier distance-cost models [9] and connects to other frameworks such as layered graph search [72] and ray-based exploration [7]. When two-dimensional search involves geometric specifications, the problem becomes more complex. Early work considered polygons [44] and disks [29], later extending to triangles, squares, and $\ell_{p}$ balls [11, 34, 52, 54]. Multi-agent extensions introduce coordination and fault-tolerance challenges [15, 20, 31], with some studies exploring worst- and average-case trade-offs [24].
In multi-agent shoreline search, spiral strategies can improve the search performance. A double spiral achieves a competitive ratio of $5.2644$ for $n=2$ [10], and for $n\geq 3$, a ray-based strategy achieves $1/\cos(\pi/n)$ with matching lower bounds [1, 40]. Partial-information variants (e.g., known slope or orientation) yield diverse bounds [57, 58, 61]. Other models consider circular targets [58], memory-limited agents on grids [41, 42, 48, 69], and navigation with cost-information trade-offs [22, 75, 73, 74].
Heterogeneous agent capabilities, such as communication [50] and speed [12], introduce further complexity. Of particular relevance to this work is heterogeneity in agent speeds, which has been rarely studied. Related results include multi-agent evacuation [66], energy-efficient sensing [77], gathering [13], patrolling [30, 62], and one-dimensional point search in bounded domains [28].
2 Definitions and Main Results
2.1 Problem Definition, Notation & Terminology
We denote by $\mathbb{R}_{*}$ the reals excluding 0, and hence by $\mathbb{R}_{*}^{2}$ the Euclidean plane excluding the origin $O=(0,0)$. A point in the plane, written in polar coordinates, is denoted by $\langle\rho,\chi\rangle$ (to differentiate from Cartesian coordinates), where $\rho$ is the radial coordinate and $\chi$ the angular coordinate (in radians). Thus, $\langle\rho,\chi\rangle=\rho(\cos\left(\chi\right),\sin\left(\chi\right))$, and in particular, $O=\langle 0,0\rangle$. We write 1 to denote a vector of all-$1$s of dimension that will be self-evident from the context. For $\psi\in[0,2\pi]$, we define the ray $\mathcal{L}_{\psi}$ as the set of positive multiples of the vector $\langle 1,\psi\rangle$, that is
$\mathcal{L}_{\psi}:=\{\langle x,\psi\rangle:x\in\mathbb{R}_{>0}\}.$
For a target point $A\in\mathbb{R}^{2}_{*}$, we denote by $\left\lVert A\right\rVert$ its Euclidean norm, that is its distance from the origin, which is always positive. We say that a point $P=\langle x,\psi\rangle\in\mathbb{R}_{*}^{2}$ exposes $A=\langle y,\psi\rangle\in\mathbb{R}_{*}^{2}$ if $x\geq y$, meaning $A$ is a convex combination of $O$ and $P$. We use the abbreviation $n]=\{0,1,2,\ldots,n-1\}$ and refer to $i\in[n]$ as mobile agent $i$ (or simply agent-$i$). By convention, agent-$n$ refers to agent-0.
We define the problem under consideration as the Multi-Speed Point Search on $\mathcal{D}$ by $n$ Agents, denoted by $\mathcal{D}\text{-}\textsc{MSP}_{n}(c)$ (or simply $\textsc{MSP}_{n}(c)$ when $\mathcal{D}$ is clear from context), where $c\in\mathbb{R}^{n}$ is the input and $\mathcal{D}\subseteq\mathbb{R}^{2}_{*}$ is the search domain. The parameter $c_{i}$ represents the speed of mobile agent $i$, with $c_{n}=c_{0}$ being the speed of agent-0. We assume speeds are given in non-increasing order, i.e., $c_{0}\geq c_{1}\geq\dots\geq c_{n-1}$, to simplify expressions in our results. Without loss of generality (after appropriate scaling), we assume $c_{0}=1$ is the highest speed. We write $\mathcal{G}_{l}$ for the geometric mean of the speeds of agents 0 through l1l-1, that is,
$\mathcal{G}_{l}:=\left(\prod\nolimits_{i\in[l]}c_{i}\right)^{1/l}.$
Observe that $c_{l}\leq\mathcal{G}_{l}$ for every $l=1,\dots,n$.
A feasible solution to $\mathcal{D}\text{-}\textsc{MSP}_{n}(c)$ consists of continuous trajectories $\{\tau_{i}\}_{i\in[n]}$ in $\mathbb{R}^{2}$, where each agent $i\in[n]$ follows a trajectory $\tau_{i}$ that depends on both $c$ and $\mathcal{D}$, satisfying:
- All agents start at the origin.
- Agent $i$ moves along $\tau_{i}$ at a speed of at most $c_{i}$.
- Each target $A\in\mathcal{D}$ is (eventually) exposed by at least one agent.
For simplicity, we say that agent $i$ exposes a point $A\in\mathcal{D}$ if some point on $\tau_{i}$ exposes $A$. For each $A\in\mathcal{D}$, let $t_{A}$ be the earliest time some agent exposes $A$. The goal is to design trajectories that minimize the search cost, defined as
$\sup_{A\in\mathcal{D}}\frac{t_{A}}{\left\lVert A\right\rVert},$
following the normalization of competitive analysis. [^1] The search cost corresponds to the notion of competitive ratio in competitive analysis. In some settings, the competitive ratio is defined with an additive term, or with an explicit restriction that excludes instances in which the reference cost becomes arbitrarily small (in our case, targets at very small distance $\left\lVert A\right\rVert$ from the origin), as a way to deal with pathological inputs. In geometric search, efficiency is instead measured directly by the ratio defining the search cost, as we do in this work. In our formulation, excluding the origin from the search domain is a sufficient and clean way to make this ratio well defined. This avoids both imposing an arbitrary positive lower bound on the target distance and introducing an additive term in the competitive ratio, which would be an unnecessary artefact in this model. Note that the origin is excluded from the search domain $\mathcal{D}$ so as to have a well-defined search cost. Intuitively, the denominator $\left\lVert A\right\rVert$ corresponds to the time required by an agent that knows the target location in advance to reach it, while the normalization yields a meaningful worst case measure by comparing the search time as a function of the distance of the hidden target from the origin, which would otherwise be unbounded.
In this work we consider the following search domains $\mathcal{D}$:
- The Plane Excluding the Origin defined as $\mathcal{P}:=\bigcup_{\psi\in[0,2\pi)}\mathcal{L}_{\psi}=\mathbb{R}^{2}_{*}$,
- the Cone of Angle $\phi$ defined as $\mathcal{C}_{\phi}:=\bigcup_{\psi\in[0,\phi)}\mathcal{L}_{\psi}$, where $\phi\in(0,\pi)$, and
- the Conic Complement of Angle $\phi$ defined as $\mathcal{W}_{\phi}:=\mathcal{P}\setminus\mathcal{C}_{\phi}$, where $\phi\in(0,\pi)$.
The notion of exposure corresponds to a well-known search/visibility criterion; a mobile agent sees a point if and only if the point lies between the agent’s location and the origin. Agents are modelled as points and do not interfere with one another; in particular, the exposure condition is defined per agent and is not affected by the locations of other agents (there is no notion of occlusion or blocking in the model). Our problem is to design trajectories for $n$ agents, considering their speed constraints, to locate a hidden target while minimizing the worst-case time required to expose a point relative to its distance from the origin.
Since trajectories are only relevant until a target is exposed, we assume each agent $i\in[n]$ moves at full speed $c_{i}$. Thus, we model agent trajectories as continuous, piecewise differentiable functions
$\tau_{i}:\mathbb{R}\to\mathbb{R}^{2},$
where each $\tau_{i}$ satisfies
$\lim_{t\to-\infty}\tau_{i}(t)=(0,0),$
with the understanding that $t$ is an abstract parameter and does not represent physical time. This convention should be read as a normalization, not as a physical requirement that the agent acts for an infinite amount of time. The parameter $t$ is a curve parameter, and physical time is obtained from arc length; saying that execution begins at $t\to$-$\infty$ simply means that the trajectory can be extended backwards so that, for every $\varepsilon>0$, the agent reaches distance at most $\varepsilon$ from the origin at some parameter value. In particular, this avoids introducing an arbitrary positive lower bound on the radius (or an arbitrary starting offset along the curve) when describing spiral-type trajectories.
We emphasize that we do not impose a common starting positive radius for the agents. For unit-speed trajectories, fixing a starting circle amounts only to a time shift and does not affect the competitive analysis, while introducing unnecessary normalization constants. More importantly, in the multi-speed setting such a requirement would be artificial: equal starting radii do not correspond to equal time or cost for agents of different speeds, and enforcing this constraint would obscure the geometric asymmetry that the algorithms in this paper are designed to exploit.
The piecewise differentiability assumption is standard in continuous motion models with bounded speed, as it ensures that instantaneous speed is well defined almost everywhere and that speed constraints can be meaningfully enforced. Curves $\tau_{i}$ are assumed to be of infinite length to guarantee that every point in the plane is exposed by some agent. That is, for each $A\in\mathcal{D}$, there exist $i\in[n]$, $t_{A}\in\mathbb{R}$, and $\lambda\in[0,1]$ such that $A=\lambda\tau_{i}(t_{A}).$
Distances are measured with respect to the standard Euclidean metric on $\mathbb{R}^{2}$, and speed bounds are interpreted with respect to arc length in this metric. Thus, the curve length from the origin to $\tau_{i}(t_{A})$ is
$\ell_{i}(t_{A}):=\int_{-\infty}^{t_{A}}\left\lVert\tau_{i}^{\prime}(t)\right\rVert\,\mathrm{d}t.$
This integral is taken over the differentiable intervals, and $\ell_{i}$ is defined for the fixed trajectory $\tau_{i}$. Since agent-$i$ is assumed to have speed $c_{i}$, the time required for agent-$i$ to reach $\tau_{i}(t_{A})$ is $\frac{\ell_{i}(t_{A})}{c_{i}},$ relating this way parameter $t_{A}$ to time for an agent of speed $c_{i}$ following trajectory $\tau_{i}(\cdot)$.
From the above terminology and notation, it follows that if $t_{A}^{i}$ is the earliest time at which $\tau_{i}(t_{A}^{i})$ exposes $A\in\mathcal{D}$ (defined as infinity if no such time exists) in $\mathcal{D}\text{-}\textsc{MSP}_{n}(c)$, then the objective of the problem is to find trajectories $\{\tau_{i}\}_{i\in[n]}$ minimizing the search cost
$\sup_{A\in\mathcal{D}}\frac{1}{\left\lVert A\right\rVert}\min_{i\in[n]}\frac{\ell_{i}(t^{i}_{A})}{c_{i}}.$
2.2 Main Contributions & Paper Organization
In this section, we summarize our main contributions, focusing on a simplified presentation of our results without explicitly describing the search algorithms. The primary focus of our work is to establish upper bounds to the search cost of $\mathcal{P}\text{-}\textsc{MSP}_{n}(c)$, i.e. the problem of searching the plane with arbitrary speed agents. As an essential component for the follow-up results, we first consider the simpler case $\mathcal{P}\text{-}\textsc{MSP}_{n}(\textbf{1})$, where all $n$ agents have unit speed. The proof of the following upper bound, based on a logarithmic-spiral type trajectory, generalizes the optimal solution for $\mathcal{P}\text{-}\textsc{MSP}_{1}(1)$ from [67]. The notation introduced in the following theorem, as well as the derived results, are essential for presenting our more general results.
Theorem 2.1
For every $n\in\mathbb{N}$, the search problem $\mathcal{P}\text{-}\textsc{MSP}_{n}(\textbf{1})$ admits a solution with search cost
$\mathcal{U}_{n}:=\tfrac{\sqrt{1+\kappa_{n}^{2}}}{\kappa_{n}}e^{2\pi\kappa_{n}/n},$
where $\kappa_{n}$ admits a closed-form expression, obtained as the root of an explicit cubic polynomial (see Lemma 4).
The proof of Theorem 2.1 is given in Section 3 and is a simple generalization of the single-agent search analysis. Some values for $\mathcal{U}_{n}$ are given in Table 2 on page 2. Quite importantly, we also show that $\mathcal{U}_{n}$ is decreasing in $n$ and approaches $1$ as $n\to\infty$.
This motivates the question of how the performance of $n$ multi-speed agents compares to that of $n$ idealized unit-speed agents in plane search. The following theorem provides one of our main technical contributions by establishing a general upper bound, expressed in terms of the previously established quantities $\mathcal{U}_{i}$ and $\kappa_{i}$. We remind the reader that $\mathcal{G}_{\ell}$ denotes the geometric mean of the speeds of the fastest $\ell$ agents $\ell$]. For convenience, we set $\mathcal{G}_{0}=1$ and define the search cost $\mathcal{U}_{0}$ for zero agents as positive infinity.
Theorem 2.2
$\mathcal{P}\text{-}\textsc{MSP}_{n}(c)$ admits a solution with search cost
$\min_{i=1,\ldots,n}\tfrac{\mathcal{U}_{i}}{\mathcal{G}_{i}}.$
In particular, this upper bound is achieved by utilizing only the agents $\ell+1]=\{0,1,\ldots,\ell\}$ to expose targets, where $\ell\in[n]$ is the largest index (corresponding to the slowest participating agent with speed $c_{\ell}$) satisfying
$c_{\ell}\geq\left(\mathcal{U}_{\ell+1}/\mathcal{U}_{\ell}\right)^{\ell+1}\mathcal{G}_{\ell}.$
The proof of Theorem 2.2 appears in Section 4, and more specifically as the corollary of a sequence of lemmata, in Section 4.4. The proof constructs logarithmic-spiral trajectories tailored to the agents’ speeds. Crucially, these trajectories introduce enough degrees of freedom to enable a competitive analysis that is agent-independent, a defining optimality condition for agents that contribute to search. Note that ordering the speeds in non-increasing order simplifies the theorem’s statement, as it maximizes the geometric mean over any subset of agents $[j]$, where $j\leq n$. Since $c_{\ell}\leq\mathcal{G}_{\ell}$ for all $\ell\in[n]$, Theorem 2.2 implies that an agent with speed $c_{\ell}$ contributes to exposing new targets only if $c_{\ell}$ is sufficiently large relative to the geometric mean of the speeds of all faster agents. This threshold is quantified by the factor
$\left(\mathcal{U}_{\ell+1}/\mathcal{U}_{\ell}\right)^{\ell+1}=1-\Theta\left(\ell^{-2/3}\right).$
These asymptotics will follow by the formula of $\mathcal{U}_{\ell}$ from Theorem 2.1 and Proposition 1 on page 1 regarding the behaviour of $\kappa_{n}$, with representative numerical values provided in Table 1. The following corollary follows immediately from Theorem 2.2.
$\ell$ 1 2 3 4 5 10 100 1000
$p_{\ell}$ 0.26668 0.33831 0.38602 0.42395 0.45579 0.56406 0.86070 0.96689
Table 1: Some indicative numerical values of $p_{\ell}:=\left(\mathcal{U}_{\ell+1}/\mathcal{U}_{\ell}\right)^{\ell+1}$ as it appears in Theorem 2.2.
Corollary 1
Searching for a point in the plane with $n$ agents of arbitrary speeds $c\in\mathbb{R}^{n}$ can be performed faster than searching with $k\leq n$ unit-speed agents, assuming that the geometric mean of the $n$ agents’ speeds is at least $\mathcal{U}_{n}/\mathcal{U}_{k},$ where $\mathcal{U}_{i}$ is the search cost for $i$ unit-speed agents, as given in Theorem 2.1.
Next, we explore alternative search trajectories for searching the plane, which, for certain agent speed ranges, strictly improve the upper bounds of Theorem 2.2 obtained from spiral-type trajectories. To this end, we establish upper bounds for searching cones and conic complements, i.e., the domains $\mathcal{C}_{\phi}$ and $\mathcal{W}_{\phi}$ for $\phi\in(0,\pi)$, which we believe are of independent interest. Our contributions focus on proposing search trajectories and analyzing their competitive performance for single unit-speed agents. Due to the technical nature of the results, the derived upper bounds are expressed in terms of the minimum values of certain functions parameterized by the angle $\phi$. Numerical evaluations of these minima, for carefully chosen $\phi$, demonstrate how to strictly improve the upper bounds of Theorem 2.2 for $\mathcal{P}\text{-}\textsc{MSP}_{2}(1,c)$ within a specific range of speeds $c\in(0,1)$.
We begin by analyzing upper bounds for searching for a point in an angle-$\phi$ cone with a unit-speed agent.
Theorem 2.3
For $\phi\in(0,\pi)$, the problem $\mathcal{C}_{\phi}\text{-}\textsc{MSP}_{1}(1)$ admits a solution with search cost
$\cos\left(\phi/2\right)\cdot\mathcal{F}_{\phi},$
where $\mathcal{F}_{\phi}$ is the minimum of the function
$\left(\frac{\sin\left(\theta\right)}{\sin\left(\theta-\phi\right)}\right)^{2}\cdot\frac{1}{\cos\left(\theta-\phi/2\right)}$
over the interval $\theta\in(\phi,\pi/2+\phi/2)$.
We prove Theorem 2.3 in Section 5. Next, we derive upper bounds for searching the conic complement of angle $\phi$ with a unit-speed agent, for $\phi\in(0,\pi)$. Importantly, the associated search cost will be strictly smaller than $\mathcal{U}_{1}$, which is optimal for searching the plane with a unit-speed agent. The following theorem establishes this result.
Theorem 2.4
For $\phi\in(0,\pi)$ and every $\lambda>1$, define
$k_{\lambda}=\frac{\ln\lambda}{2\pi}+\frac{\lambda-1}{\sqrt{2\lambda(1-\cos\left(\phi\right))}}\left(1-\frac{\phi}{2\pi}\right).$
Then, the problem $\mathcal{W}_{\phi}\text{-}\textsc{MSP}_{1}(1)$ admits a solution with search cost $\mathcal{R}_{\phi}$, where $\mathcal{R}_{\phi}$ is the minimum of the function
$\frac{\sqrt{\lambda^{2}-2\lambda\cos(\phi)+1}}{\lambda-1}e^{2\pi k_{\lambda}}$
in the interval $\lambda\in(1,\infty)$. Moreover, for every $\phi\in(0,\pi)$ this value is strictly smaller than $\mathcal{U}_{1}$.
We prove Theorem 2.4 in Section 6. Having established these bounds, we now show that spiral-type algorithms may not be optimal for searching the plane with multiple agents. We focus on $\mathcal{P}\text{-}\textsc{MSP}_{2}(1,c)$, where two agents, one unit-speed and one speed-$c$, search the plane. We derive an alternative upper bound to Theorem 2.2, achieved through a hybrid strategy that combines cone and conic complement search techniques.
Theorem 2.5
Let
$\gamma(\phi)=\cos\left(\phi/2\right)\frac{\mathcal{F}_{\phi}}{\mathcal{R}_{\phi}},$
where $\mathcal{F}_{\phi}$ and $\mathcal{R}_{\phi}$ are as defined in Theorems 2.3 and 2.4, respectively. Then, for every $\phi\in(0,\pi)$, the problem $\mathcal{P}\text{-}\textsc{MSP}_{2}(1,\gamma(\phi))$ admits a solution with search cost equal to $\mathcal{R}_{\phi}$, or also equal to
$\cos\left(\phi/2\right)\frac{\mathcal{F}_{\phi}}{\gamma(\phi)}.$
We prove Theorem 2.5 in Section 7. At a high level, we use Theorem 2.4 to assign the unit-speed agent to search a conic complement, ensuring a search cost strictly smaller than $\mathcal{U}_{1}$. We then assign the speed-$c$ agent to search the corresponding cone using the trajectory from Theorem 2.3, carefully tuning the cone angle $\phi=\phi(c)$ so that its search cost does not exceed that of the faster agent. When the cone is sufficiently narrow, i.e. $\phi$ is small enough, even a slow agent can explore it efficiently. The relationship between the cone angle $\phi$ and the required agent speed is captured by the function $\gamma(\phi)$, which defines the minimum speed needed to ensure balanced competitive costs.
Notably, from our analysis, Theorem 2.2 guarantees an upper bound of $\mathcal{U}_{1}$ for
$c\leq(\mathcal{U}_{2}/\mathcal{U}_{1})^{2}\approx 0.266687$
when searching the plane with two agents (see also Table 1), one unit-speed and one with speed $c$. This matches the bound achieved by a single unit-speed agent, meaning the speed-$c$ agent does not contribute to exposing any points in this range. Furthermore, we show that if
$c\leq 1/\mathcal{U}_{1}\approx 0.0578391,$
then $\mathcal{U}_{1}$ is the optimal search cost for $\mathcal{P}\text{-}\textsc{MSP}_{2}(1,c)$. Theorem 2.5 improves the search cost of $\mathcal{U}_{1}$, when $c>1/\mathcal{U}_{1}$ and for speeds slightly exceeding $0.266687$. This threshold corresponds to the speed at which both agents, with speeds $1$ and $c$, expose new targets in the spiral-type trajectory used in Theorem 2.2. Our numerical results, summarized in the next theorem, show that our spiral-type algorithm, based on non-intersecting trajectories and inducing an agent-independent competitive ratio (a defining optimality condition), is provably suboptimal for low-speed agents. This provides strong evidence that spiral-type trajectories may be suboptimal more generally for the planar search problem with arbitrary-speed agents. Since Theorem 2.2 already yields strong upper bounds for larger values of $c$, we focus below on the case of small speeds.
Theorem 2.6
For $c\in(0,0.5]$, the problem $\mathcal{P}\text{-}\textsc{MSP}_{2}(1,c)$ admits the upper bound depicted in Figure 1. Moreover, for $c\geq 0.268872$, the bound is achieved by a logarithmic-spiral-type algorithm, as in Theorem 2.2. For $0.0578391<c<0.268872$, the upper bound is strictly better than $\mathcal{U}_{1}$, and is achieved by a strategy that partitions the plane into a cone $\mathcal{C}_{\phi}$ (searched by the speed-$c$ agent) and a wedge $\mathcal{W}_{\phi}$ (searched by the unit-speed agent), for an appropriate choice of $\phi\in(0,\pi)$.
Figure 1: The orange curve represents the upper bound from Theorem 2.2, achieved by the Off-Set Algorithm 2 for searching the plane with agent speeds $1,c$, as a function of $c\in(0,1)$. The transition value where the search cost is not differentiable in the speed is $c=(\mathcal{U}_{2}/\mathcal{U}_{1})^{2}\approx 0.266687$, and note that past this threshold speed value, the induced search cost is strictly lower than $\mathcal{U}_{1}$. The blue curve represents the upper bound from Theorem 2.5, achieved by the Cone-Wedge Hybrid Algorithm 6 for searching the plane with agent speeds $1,c$. The right panel zooms in on the region where the two bounds converge, showing that the latter strategy is strictly better for searching the plane even for $c\in(0.266687,0.268872)$ in which both the speed-$1$ and speed-$c$ agents expose new targets following the Off-Set Algorithm 2.
The improvement achieved by the hybrid strategy occurs only in a narrow interval of speed ratios. For very small values of $c$, the slow agent cannot contribute meaningfully, and the bound coincides with the unit-speed spiral strategy; for larger values of $c$, the spiral-based bound $\mathcal{U}_{2}/\sqrt{c}$ dominates. The intermediate regime visible in Figure 1 arises from a numerical optimization of the cone angle in the hybrid construction, and reflects a geometric trade-off between restricting the slow agent to a narrow cone and increasing the cost of searching the complementary region. We prove Theorem 2.6 in Section 7.
3 Searching the Plane with $n$ Unit-Speed Agents
In this section, we prove Theorem 2.1, establishing upper bounds for the problem $\mathcal{P}\text{-}\textsc{MSP}_{n}(\textbf{1})$, where $n$ unit-speed agents search the plane. The derived upper bounds naturally generalize the provably optimal algorithm for $\textsc{MSP}_{1}(1)$ due to [67]. The optimal trajectory for this problem follows a logarithmic spiral, parameterized by the expansion factor $k>0$. The trajectories for $n$ unit-speed agents are based on the same principle and are defined as follows.
Figure 2: The depiction of trajectories of Algorithm 1. In this example, we use $k=0.356$, $n=30$, and we show the trajectories of agents $0,1,2,29$, along with their positions when $t=t_{s}=2\pi s$, for some $s\in\mathbb{Z}$. All agents lie on the circle of radius $e^{kt_{s}}$, while agent-0 lies on ray $\mathcal{L}_{0}$. The last agent who previously hit $\mathcal{L}_{0}$ was agent-(n1)(n-1).
Algorithm 1 Uniform Spiral Trajectory for $\mathcal{P}\text{-}\textsc{MSP}_{n}(\textbf{1})$
0:$k>0$
Output Trajectories: $f_{i}^{k}(t)=\langle e^{kt},t-\tfrac{2\pi}{n}i\rangle,\quad t\in\mathbb{R},\quad i\in[n].$
The logarithmic spiral guarantees that the distance of newly exposed points from the origin grows proportionally to the length of the trajectory. This property is necessary and sufficient for achieving a constant search cost. Using the same expansion factor across agents further guarantees that trajectories do not overlap.
Algorithm 1 is depicted in Figure 2. The execution of the algorithm begins at $t\to$-$\infty$. The condition $k>0$ is necessary for correctness, ensuring that every point in the plane is eventually exposed. For any fixed parameter value $t$, all agents have the same distance $e^{kt}$ from the origin, while their angular separation is exactly $2\pi/n$. Consequently, at every snapshot of the algorithm, all agents lie on a common circle and their positions form a regular $n$-gon (up to rotation and scaling as $t$ varies).
The following lemma establishes that Algorithm 1 is indeed a feasible search strategy.
Lemma 1
For every $k>0$ and $n\in\mathbb{N}$, Algorithm 1 is feasible for $\mathcal{P}\text{-}\textsc{MSP}_{n}(\textbf{1})$.
Proof
For every $i\in[n]$ and $k>0$, we observe that
$\lim_{t\rightarrow-\infty}\left\lVert f_{i}^{k}(t)\right\rVert=\lim_{t\rightarrow-\infty}e^{kt}=0,$
which confirms that agents start from the origin.
Next, we show that every point $A=\langle x,\psi\rangle$, with $x>0$ and $\psi\in[0,2\pi]$, is exposed. Consider the smallest $t$ such that $e^{kt}\geq x$, which exists since $k>0$. Define the smallest $s\in\mathbb{Z}$ such that $2\pi s+\psi\geq t$. Then, $f_{0}^{k}(2\pi s+\psi)\in\mathcal{L}_{\psi},$ and $\left\lVert f_{0}^{k}(2\pi s+\psi)\right\rVert=e^{2\pi s+\psi}\geq e^{kt}\geq x,$ meaning that agent-0 exposes $A$. ∎
The parameter $t$ in Algorithm 1 is related to, but not equal to, time when unit-speed agents follow trajectories $f_{i}^{k}(t)$. The next lemma explicitly describes this relationship and shows that the trajectory lengths $f_{i}^{k}(t)$ are invariant with respect to $i$.
Lemma 2
Fix $\phi\in\mathbb{R}$, $\alpha,\beta\in\mathbb{R}$ with $\alpha>0$, and consider $h:\mathbb{R}\mapsto\mathbb{R}^{2}$ with $h(t)=\langle e^{\alpha t+\beta},t+\phi\rangle$. Then
$\int_{-\infty}^{y}\left\lVert h^{\prime}(t)\right\rVert dt=\frac{\sqrt{1+\alpha^{2}}}{\alpha}e^{\alpha y+\beta}.$
Proof
In Cartesian coordinates, we have $h(t)=e^{\alpha t+\beta}\left(\cos\left(t+\phi\right),\sin\left(t+\phi\right)\right)$. Therefore,
$\mathrm{d}h(t)/\,\mathrm{d}t=e^{\alpha t+\beta}\left(\alpha\cos(t+\phi)-\sin(t+\phi),\alpha\sin(t+\phi)+\cos(t+\phi)\right),$
and hence
$\left\lVert\,\mathrm{d}h(t)/\,\mathrm{d}t\right\rVert^{2}$ $=e^{2(\alpha t+\beta)}\left(\left(\alpha\cos(t+\phi)-\sin(t+\phi)\right)^{2}+\left(\alpha\sin(t+\phi)+\cos(t+\phi)\right)^{2}\right)$
$=e^{2(\alpha t+\beta)}\left(1+\alpha^{2}\right).$
Therefore,
$\int_{-\infty}^{y}\left\lVert\,\mathrm{d}h(t)/\,\mathrm{d}t\right\rVert\,\mathrm{d}t=\sqrt{1+\alpha^{2}}\int_{-\infty}^{y}e^{\alpha t+\beta}\,\mathrm{d}t=\frac{\sqrt{1+\alpha^{2}}}{\alpha}e^{\alpha y+\beta},$
where the last equality is due to the fact that $\alpha>0$. ∎
As an immediate corollary, we have that for the trajectories of Algorithm 1, and for each $i\in[n]$, $\int_{-\infty}^{y}\left\lVert f_{i}^{\prime}(t)\right\rVert dt=\frac{\sqrt{1+k^{2}}}{k}e^{ky}.$ We can now derive the following upper bound, which for $n=1$, and a proper choice of $k$, gives the known optimal result for $\mathcal{P}\text{-}\textsc{MSP}_{1}(1)$ due to [67].
Lemma 3
The search cost of Algorithm 1 and input $k>0$ for $\mathcal{P}\text{-}\textsc{MSP}_{n}(\textbf{1})$ equals
$\frac{\sqrt{1+k^{2}}}{k}e^{2\pi k/n}.$
Proof
By the symmetry of the trajectories $\{f_{i}^{k}(t)\}_{i\in[n]}$, defined for all $t\in\mathbb{R}$, and by appropriately rotating the plane about the origin, we may assume that the target lies on ray $\mathcal{L}_{0}$. Furthermore, due to the symmetry of all trajectories, we may assume that the target is exposed by agent-0.
Define $t_{s}:=2\pi s$ for $s\in\mathbb{Z}$. That is, the target $A_{x}=(x,0)$ is first exposed when agent-0 reaches $f_{0}^{k}(t_{s})$ for some $s\in\mathbb{Z}$. It follows that $x\leq e^{kt_{s}}=e^{k(2\pi s)}$. Moreover, by Lemma 2, agent-0 reaches this point at time
$\ell_{0}(t_{s})=\frac{\sqrt{1+k^{2}}}{k}e^{kt_{s}}=\frac{\sqrt{1+k^{2}}}{k}e^{k(2\pi s)}.$
Next, we establish a lower bound on $x$ under the assumption that agent-0 is the first to expose $A_{x}$. Since time increases with the parameter $t$ along the trajectories, the largest $t<t_{s}$ for which $f_{j}^{k}(t)$ was on the positive horizontal axis is given by
$t+\frac{2\pi(n-j)}{n}\leq t_{s}=2\pi s,$
which implies
$t\leq 2\pi s-\frac{2\pi(n-j)}{n}\leq 2\pi s-\frac{2\pi}{n},$
where the last inequality is tight for j=n1j=n-1, see also Figure 2. Thus, before $A_{x}$ was exposed, the last agent to cross the positive horizontal axis was agent-(n1)(n-1), reaching the point $f_{n-1}^{k}(2\pi s-2\pi/n)$ and exposing all points in the closed interval $\left[0,e^{k(2\pi s-2\pi/n)}\right]$. It follows that $x>e^{k(2\pi s-2\pi/n)},$ with strict inequality. Hence, the worst-case search cost for points exposed by agent-0 is
$\sup_{s\in\mathbb{Z}}\sup_{x\in\left(e^{k(2\pi s-2\pi/n)},e^{k(2\pi s)}\right]}\frac{\ell_{0}(t_{s})}{x}=\sup_{s\in\mathbb{Z}}\frac{\sqrt{1+k^{2}}}{k}\frac{e^{k(2\pi s)}}{e^{k(2\pi s-2\pi/n)}}=\frac{\sqrt{1+k^{2}}}{k}e^{2\pi k/n}.$
In particular, these calculations show that the search cost is independent of $s$ in the above analysis and also independent of agent-$i$, as established by the symmetry argument at the beginning of the proof. ∎
Minimizing the search cost expression in Lemma 3 with respect to $k=k(n)$ yields the following upper bounds for $\mathcal{P}\text{-}\textsc{MSP}_{n}(\textbf{1})$, along with the corresponding optimal expansion factors, denoted by $\kappa_{n}$. The next lemma provides a closed formula for $\kappa_{n}$.
Lemma 4
For each $n\in\mathbb{N}$, the search cost of trajectories $\{f_{i}^{k}(t)\}_{i\in[n]}$ for $\mathcal{P}\text{-}\textsc{MSP}_{n}(\textbf{1})$, as seen in Lemma 3, is convex in expansion factor $k$. Moreover, the corresponding optimizers $k=\kappa_{n}$ are given as the unique positive root to $k^{3}+k=n/2\pi$, and therefore
$\kappa_{n}:=\sqrt[3]{\sqrt{\frac{n^{2}}{16\pi^{2}}+\frac{1}{27}}+\frac{n}{4\pi}}-\sqrt[3]{\sqrt{\frac{n^{2}}{16\pi^{2}}+\frac{1}{27}}-\frac{n}{4\pi}}.$
Proof
Consider the function $z(k)=\frac{\sqrt{1+k^{2}}}{k}e^{k/a}$ parameterized by $a\in\mathbb{R}_{>0}$. We compute
$\frac{\,\mathrm{d}}{\,\mathrm{d}k}z(k)=\frac{e^{k/a}\left(-a+k^{3}+k\right)}{ak^{2}\sqrt{k^{2}+1}},$
so that any $k_{a}$ satisfying $k_{a}^{3}+k_{a}=a$ is a critical point. We show that this corresponds to a global minimum. Indeed, computing the second derivative, we obtain
$\tfrac{\,\mathrm{d}^{2}}{\,\mathrm{d}k^{2}}z(k)=e^{k/a}\frac{a^{2}\left(3k^{2}+2\right)+\left(k^{3}+k\right)\left(k^{3}+k-2a\right)}{a^{2}k^{3}\left(k^{2}+1\right)^{3/2}}.$
It follows that
$\frac{\,\mathrm{d}^{2}}{\,\mathrm{d}k^{2}}z(k_{a})=\frac{\left(3k_{a}^{2}+1\right)e^{k_{a}/a}}{k_{a}^{3}\left(k_{a}^{2}+1\right)^{3/2}}>0,$
confirming that $k_{a}$ is indeed a minimizer.
More generally, we analyze the discriminant of the numerator in $\tfrac{\,\mathrm{d}^{2}}{\,\mathrm{d}k^{2}}z(k)$, which equals
$a^{2}\left(3k^{2}+2\right)+\left(k^{3}+k\right)\left(k^{3}+k-2a\right),$
and we view it as a degree-2 polynomial in $a$. The discriminant evaluates to
$k^{2}-5k^{4}-7k^{6}-3k^{8}<0.$
Since this is negative, the polynomial maintains a constant sign, which is determined by the leading coefficient, $3k^{2}+2>0$. It follows that $\frac{\,\mathrm{d}^{2}}{\,\mathrm{d}k^{2}}z(k)>0$ for all $k$, proving that $z(k)$ is convex.
Next, we derive a closed-form expression for $k_{a}$, which satisfies the equation $k^{3}+k-a=0.$ The discriminant of this degree-3 polynomial is $4(1)^{3}-27(-a)^{2}=-4-27a^{2}<0.$ Since the discriminant is negative, the equation has only one real root, given by Cardano’s formula, and that is
$k_{a}=\left(\sqrt{\frac{a^{2}}{4}+\frac{1}{27}}+\frac{a}{2}\right)^{\frac{1}{3}}-\left(\sqrt{\frac{a^{2}}{4}+\frac{1}{27}}-\frac{a}{2}\right)^{\frac{1}{3}}.$
Setting $\kappa_{n}=k_{a}$ with $a=n/2\pi$ completes the proof. ∎
The proof of Theorem 2.1 is obtained now as an easy corollary.
Proof(of Theorem 2.1)
We use the trajectories from Algorithm 1 with expansion factor $\kappa_{n}$ from Lemma 4. By Lemma 3, the resulting search cost is as stated in the theorem. ∎
$n$ 1 2 3 4 5 6 7
Search Cost $\mathcal{U}_{n}$ 17.28935 8.92852 6.22135 4.90385 4.13049 3.62351 3.2659
$\kappa_{n}$ 0.155402 0.293124 0.409031 0.506602 0.590194 0.663214 0.728099
Table 2: Search cost for $\textsc{MSP}_{n}(\textbf{1})$ and corresponding optimizer expansion factors $k=\kappa_{n}$ values for trajectories $\{f_{i}^{k}\}_{i\in[n]}$, when $n=1,\ldots,7$.
Table 2 provides some indicative numerical values for search costs $\mathcal{U}_{n}$, for small values of $n$. The values $\mathcal{U}_{1}$ and $\kappa_{1}$ are the known values that have given the provably optimal algorithm to $\mathcal{P}\text{-}\textsc{MSP}_{1}(1)$ due to [67]. Below, we explore the behavior of $\mathcal{U}_{n},\kappa_{n}$ as a function of $n$, which will be useful later.
Proposition 1
Expression $\mathcal{U}_{n}$ of Theorem 2.1 is strictly decreasing in $n$, and $\lim_{n\rightarrow\infty}\mathcal{U}_{n}=1$. Moreover, the optimizer expansion factor $\kappa_{n}$ is positive, it is strictly increasing, and strictly concave in $n$, with $\kappa_{n}=\Theta\left(n^{1/3}\right)$.
Proof
By construction, $\kappa_{n}$ is the unique positive solution to $q(x):=x^{3}+x=\frac{n}{2\pi}.$ Function $q(x)$ is strictly increasing and strictly convex on $x>0$, therefore it has an inverse. Hence, $\kappa_{n}$ is given as $q^{-1}(n/2\pi)$, and as a result, $\kappa_{n}$ is positive, it is strictly increasing, and strictly concave in $n$. Among others, this also implies that $\kappa_{n}^{\prime}>0$.
Next, we calculate
$\frac{\,\mathrm{d}}{\,\mathrm{d}a}k_{a}=\frac{\sqrt[3]{\sqrt{81a^{2}+12}-9a}+\sqrt[3]{\sqrt{81a^{2}+12}+9a}}{\sqrt[3]{2}\sqrt[6]{3}\sqrt{27a^{2}+4}}.$
and we see that
$\frac{\,\mathrm{d}}{\,\mathrm{d}n}\mathcal{U}_{n}=-\frac{e^{\frac{2\pi\kappa_{n}}{n}}\left(n^{2}\kappa_{n}^{\prime}+2\pi\kappa_{n}\left(\kappa_{n}^{2}+1\right)\left(\kappa_{n}-n\kappa_{n}^{\prime}\right)\right)}{n^{2}\kappa_{n}^{2}\sqrt{\kappa_{n}^{2}+1}},$
Therefore, $\mathcal{U}_{n}$ is decreasing as long as we show that $\kappa_{n}-n\kappa_{n}^{\prime}>0$. This is what we show next. Again, because $a=n/2\pi$, we see that
$\frac{\kappa_{n}}{n\tfrac{\,\mathrm{d}}{\,\mathrm{d}n}\kappa_{n}^{\prime}}=\frac{k_{a}}{a\tfrac{\,\mathrm{d}}{\,\mathrm{d}a}k_{a}^{\prime}}=\frac{(\alpha-\beta)\left((\alpha-\beta)^{2}+5\cdot 12^{1/3}\right)}{6a}$ (1)
where $\alpha=(9a+\sqrt{12+81a^{2}})^{1/3}$ and $\beta=(-9a+\sqrt{12+81a^{2}})^{1/3}$. Now observe that $\alpha>9a$ for all $a>0$, and therefore $\alpha>\beta$, concluding that expression (1) is indeed positive, which in return, implies that $\mathcal{U}_{n}$ is decreasing in $n$.
Lastly, we show that $\mathcal{U}_{n}$ tends to $1$, as the number of agents $n$ grows. For that, we see from Lemma 4 that $\kappa_{n}=\Theta\left(n^{1/3}\right)$, and hence $\kappa_{n}/n=\Theta\left(n^{-2/3}\right)$. But then,
$\lim_{n\rightarrow\infty}\mathcal{U}_{n}=\lim_{n\rightarrow\infty}\frac{\sqrt{1+\kappa_{n}^{2}}}{\kappa_{n}}e^{2\pi\kappa_{n}/n}=\left(\lim_{n\rightarrow\infty}\frac{\sqrt{1+\kappa_{n}^{2}}}{\kappa_{n}}\right)e^{\lim_{n\rightarrow\infty}2\pi\kappa_{n}/n}=1\cdot 1=1.$
4 Searching the Plane with $n$ Agents of Arbitrary Speeds
In this section, we prove Theorem 2.2, providing upper bounds for the problem $\mathcal{P}\text{-}\textsc{MSP}_{n}(c)$, where $n$ agents of arbitrary speeds search the plane. First, in Section 4.1, we introduce the Off-Set Spiral Trajectory, a general algorithm for the problem, presented as Algorithm 2. Then, in Section 4.2, we analyze its general properties with respect to its parameters and the input speeds $c\in\mathbb{R}^{n}$. Building on these observations, Section 4.3 determines an appropriate subset of agents and selects optimal parameters for running the Off-Set algorithm. This leads to Theorem 4.1 (page 4.1), which establishes certain conditions on the input speeds and upper bounds for $\mathcal{P}\text{-}\textsc{MSP}_{n}(c)$. The parameters used in the theorem are computed by Algorithm 3 and serve as input to Algorithm 2. Finally, in Section 4.4, we simplify the premises of Theorem 4.1, leading to the proof of Theorem 2.2.
Our construction should be viewed as a best-possible extension of spiral-based search to heterogeneous speeds, under the strong requirement that all participating agents incur the same competitive cost. One of the goals of the paper is to understand whether this natural generalization remains optimal once agent speeds differ.
4.1 The Speed-Dependent Off-Set Spiral Algorithm for $\mathcal{P}\text{-}\textsc{MSP}_{n}(c)$
In this section, we generalize the trajectory $\{f_{i}^{k}\}_{i\in[n]}$ for $\textsc{MSP}_{n}(\textbf{1})$ in order to propose a polynomial-time algorithm that solves $\textsc{MSP}_{n}(c)$ for arbitrary $c\in\mathbb{R}^{n}$. Recall that we made the assumptions about $\{c_{i}\}_{i\in[n]}$ are given in non-increasing order and that $c_{0}=1$. As a reminder, we also refer to agent-0 as agent-$n$, and we use $c_{0}=c_{n}=1$. In that direction, we propose a family of spiral-type search algorithms, as described in Algorithm 2, which we call Off-Set Trajectories, in which agents’ trajectories do not intersect. The key difference from Algorithm 1 is that agents in the new trajectories are assigned arbitrary angular separations, parameterized by values $\phi_{i}$, which will later be defined in terms of the agents’ speeds. These degrees of freedom, in addition to the spiral expansion parameter $k$, enable a competitive analysis that is agent-independent, a defining optimality property, thereby improving the overall efficiency of the agents that will contribute to search.
Algorithm 2 Off-Set Trajectory for $\mathcal{P}\text{-}\textsc{MSP}_{n}(c)$
[…]
Output Trajectories: $g_{i}^{k}(t)=\langle c_{i}\cdot e^{kt},t+\phi_{i}\rangle,\quad t\in\mathbb{R},\quad i\in[n].$
For any fixed parameter value $t$, the distance of agent $i$ from the origin equals $c_{i}e^{kt}$, and thus the distances of any two agents remain proportional. Consequently, as $t$ varies, the point configuration $\{g_{i}^{k}(t)\}_{i\in[n]}$ evolves by uniform scaling and rotation, and any two such configurations are similar. This contrasts with the unit speed case, where all agents have the same distance from the origin at each snapshot, and the configuration forms a regular $n$-gon. Thus, in the multispeed setting, the configuration need not be regular, but its shape is preserved over time up to similarity, see Figure 3.
Figure 3: The depiction of trajectories of Algorithm 2. In this example, we use $n=3$ agents with speeds $c_{0}=1$, $c_{1}=0.9$, and $c_{2}=0.7$, and we plot the corresponding offset logarithmic spirals (as defined by the algorithm). We also show two snapshots of the agents, at two distinct parameter values, together with the triangles formed by connecting the three agent positions in each snapshot. In contrast to the unit speed case (Figure 2), where all agents lie on the same circle at every snapshot and hence form a regular 33-gon, the multispeed agents lie at different distances from the origin; these distances remain proportional, and the two triangles are similar (their side lengths scale by the same factor between snapshots).
Note that the parameters $\phi_{0}=2\pi$ and $\phi_{n}=0$ are fixed and correspond to agent-0 (also referred to as agent-$n$ for notational convenience). Due to the periodicity of trigonometric functions, we have $g_{0}(t)=g_{n}(t)$, ensuring cyclic behavior in the agents’ trajectories. As before, the condition $k>0$ is necessary for the correctness of the search algorithm, guaranteeing that every point in the plane is eventually exposed. The ordering $\phi_{0}\geq\phi_{1}\geq\dots\geq\phi_{n-1}\geq 0$ is a normalization that determines the appropriate sequence in which agents should be considered based on their speeds. This ordering ensures that whenever agent $i$ reaches a given ray, the next agent to do so is agent $(i+1)\bmod n$. Consequently, every ray is visited in cyclic order by agents $0,1,2,\dots,n-1$, repeating indefinitely.
The main contribution of this section is a polynomial-time algorithm that takes as input $\{c_{i}\}_{i\in[n]}$ and determines parameters for the Off-Set Algorithm 2, yielding strong upper bounds for $\mathcal{P}\text{-}\textsc{MSP}_{n}(c)$. Notably, we show that for an arbitrary input $c\in\mathbb{R}^{n}$, not all agents necessarily expose new points in a solution. Thus, minimizing the search cost of the Off-Set algorithm for $\textsc{MSP}_{n}(c)$ reduces to identifying a subset $A\subseteq[n]$ and parameters $k>0$ and $\{\phi_{i}\}_{i\in A}$ that minimize the search cost.
To compute these parameters, we propose Algorithm 3, which utilizes precomputed expansion factors $\kappa_{j}$ from Lemma 4 and the search cost $\mathcal{U}_{j}$ for $j$ unit-speed agents from Theorem 2.1. Algorithm 3 returns parameters $\phi_{j}$ only for the agents that actively contribute to the search. Specifically, it identifies a key agent and all agents at least as fast, while ignoring slower agents, as they do not expose new points. The values of $\phi_{j}$ for these ignored agents are irrelevant to the solution.
Algorithm 3 Off-Set Trajectory Parameter Selection
[…]
1:  Find index $l\in[n]$ that satisfies $c_{l}\geq e^{\frac{-2\pi\kappa_{l+1}}{l+1}}\mathcal{G}_{l+1}$ (or equivalently $c_{l}\geq e^{\frac{-2\pi\kappa_{l+1}}{l}}\mathcal{G}_{l}$).
2:  Set $\mu\leftarrow\operatorname*{arg\,min}_{m\in[l]}\mathcal{U}_{m+1}/\mathcal{G}_{m+1}$ and $\nu\leftarrow\mu+1$.
3:Output: $\mu,\nu$, and $\phi_{j}=\frac{2\pi(\nu-j)}{\nu}-\frac{j}{\kappa_{\nu}}\ln\left(\mathcal{G}_{j}/\mathcal{G}_{\nu}\right)$, for $j\in[\mu]$.
We emphasize that the index $l$ in Step 1 of Algorithm 3 can be chosen as the largest index satisfying the condition, corresponding to the slowest agent. This choice will later optimize our upper bounds. The goal in subsequent Sections 4.2 and 4.3 is to prove the next theorem, which can be seen as precursor to Theorem 2.2.
Theorem 4.1
Let $\mu$ and $\nu=\mu+1$ be the indices computed by Algorithm 3 on input $c\in\mathbb{R}^{n}$, along with parameters $\{\phi_{i}\}_{i\in[\mu]}$. Then, we have $\mu\geq 0$, which implies $\nu\geq 1$, and the parameters $\{\phi_{i}\}_{i\in[\mu]}$ along with expansion factor $k=\kappa_{\nu}$ are valid for the Off-Set Algorithm 2 applied to the problem $\mathcal{P}\textrm{-}\textsc{MSP}_{n}(c)$. Moreover, the induced search cost is $\mathcal{U}_{\nu}/\mathcal{G}_{\nu}$, where the value $\mathcal{U}_{\nu}$ is the search cost for searching with $\nu$ unit-speed agents, and is given in Theorem 2.1.
4.2 Properties of the Off-Set Trajectories
In this section, we identify and analyze key properties of Algorithm 2. Notably, in this section we do not assume that the agent speeds $c_{i}$ are given in non-increasing order. We begin by showing that agents following the trajectories $\{g_{i}^{k}(t)\}_{i\in[n]}$ search the plane in an invariant pattern.
Lemma 5
For the Off-Set search strategy $\{g^{k}_{i}\}_{i\in[n]}$ to $\textsc{MSP}_{n}(c)$, and for every $i\in[n]$, we have $\frac{1}{c_{i}}\ell_{i}(t_{0})=\ell_{0}(t_{0})$, for all $t_{0}\in\mathbb{R}$. As a result, at any moment of the execution of the search strategy, any two agents maintain the same angular distance.
Proof
Fix some $t_{0}\in\mathbb{R}$, and agent $i$ with speed $c_{i}$. We observe that the angular distance between points $g_{0}^{k}(t_{0})$ and $g_{i}^{k}(t_{0})$ equals $2\pi-\phi_{i}$. Hence, the two agents preserve their angular distance if and only if they reach the corresponding points $g_{0}^{k}(t_{0})$ and $g_{i}^{k}(t_{0})$ at the same time. We argue next that the length of trajectory $g^{k}_{i}$ up to $t=t_{0}$ is $\ell_{i}(t_{0})=c_{i}\ell_{0}(t_{0})$. Indeed, let $h:\mathbb{R}\to\mathbb{R}^{2}$ be as in the statement of Lemma 2. We note that $h(t)=g_{0}^{k}(t)$, since $\phi_{0}=2\pi$ and $c_{0}=1$. But then, by the same lemma, and for any speed-$c_{i}$, we have
$\ell_{i}(t_{0})=\int_{-\infty}^{t_{0}}\left\lVert\,\mathrm{d}g_{i}^{k}(t)/\,\mathrm{d}t\right\rVert\,\mathrm{d}t=c_{i}\int_{-\infty}^{t_{0}}\left\lVert\,\mathrm{d}h(t)/\,\mathrm{d}t\right\rVert\,\mathrm{d}t=c_{i}\int_{-\infty}^{t_{0}}\left\lVert\,\mathrm{d}g_{0}^{k}(t)/\,\mathrm{d}t\right\rVert\,\mathrm{d}t=c_{i}\ell_{0}(t_{0}),$
as promised. But then, the speed-$c_{i}$ agent reaches point $g_{i}^{k}(t_{0})$ in time $\frac{1}{c_{i}}\ell_{i}(t_{0})=\frac{1}{c_{i}}c_{i}\ell_{0}(t_{0})=\ell_{0}(t_{0})$, which is the same as the time that the unit-speed agent needs to reach point $g_{0}^{k}(t_{0})$. That is, the angular distance between agent-$i$ and agent-0 is preserved, hence the angular distance between any two agents is preserved. ∎
The next lemma establishes a condition for every agent to expose new points, and recall that for the Off-Set trajectory parameters, we assume that $2\pi=\phi_{0}\geq\phi_{1}\geq\ldots\geq\phi_{n-1}\geq 0$.
Lemma 6
For the Off-Set search strategy $\{g^{k}_{i}\}_{i\in[n]}$ to $\textsc{MSP}_{n}(c)$ with parameters $\{\phi_{i}\}_{i\in[n]}$ and $k>0$, all agents expose new points if and only if
$c_{i}e^{-k\phi_{i}}<c_{i+1}e^{-k\phi_{i+1}},i\in[n].$
Here, agent $n$ is identified as agent 0 and $c_{0}=1$, $\phi_{n}=0$. Moreover, if the above condition holds, then for every $s\in\mathbb{Z}$ and for $i\in[n]$, agent $i+1$ is the first to expose all points $(x,0)\in\mathbb{R}^{2}$ on the ray $\mathcal{L}_{0}$, where
$x\in\left(c_{i}e^{2\pi ks-k\phi_{i}},c_{i+1}e^{2\pi ks-k\phi_{i+1}}\right].$
Proof
Consider $i\in[n]$ and agents $i$ and $i+1$ with speeds $c_{i}$ and $c_{i+1}$, respectively, so that $\phi_{i}\geq\phi_{i+1}$. Let $t_{0}$ be the time at which agent-$i$ reaches $\mathcal{L}_{0}$. By rotating the plane around the origin, we may assume that $t_{0}+\phi_{i}=2\pi s$ for some $s\in\mathbb{Z}$. Using Lemma 5, we see that by time $\ell_{0}(t_{0})$, all points in the closed interval $0,c_{i}e^{kt_{0}}]=[0,c_{i}e^{2\pi sk-k\phi_{i}}]$ are already exposed by agent-$i$. Since time increases with $t_{0}$ and $\phi_{i}\geq\phi_{i+1}$, it follows that $0\leq\phi_{i}-\phi_{i+1}\leq 2\pi.$ The next time agent-$(i+1)$ reaches the positive horizontal axis is when it arrives at
$g_{i+1}^{k}(t_{0}+\phi_{i}-\phi_{i+1})=g_{i+1}^{k}(2\pi s-\phi_{i+1})=c_{i+1}e^{2\pi ks-k\phi_{i+1}}(1,0).$
New points are exposed precisely when
$c_{i+1}e^{2\pi ks-k\phi_{i+1}}>c_{i}e^{2\pi ks-k\phi_{i}},$
or equivalently,
$c_{i+1}e^{-k\phi_{i+1}}>c_{i}e^{-k\phi_{i}}.$
Furthermore, the newly exposed points lie in the semi-open (and semi-closed) interval as stated in the lemma. ∎
We now calculate the search cost of the Off-Set algorithm under specific conditions on its parameters and input speeds. To do so, we introduce the agent-$i$ search cost, defined as the search cost of the Off-Set algorithm restricted to points exposed by agent-$i$. If only a subset of agents $A\subseteq[n]$ expose new points, the overall search cost of the algorithm is given by the maximum agent-$i$ search cost over all $i\in A$.
Lemma 7
Consider the Off-Set search strategy $\{g^{k}_{i}\}_{i\in[n]}$ with parameters $\{\phi_{i}\}_{i\in[n]}$ and expansion factor $k>0$, and assume that all agents expose new points. Then, the agent-$(i+1)$ search cost equals
$\frac{\sqrt{1+k^{2}}}{k}\frac{1}{c_{i}}e^{-k(\phi_{i+1}-\phi_{i})},~~i\in[n],$
where agent-$n$ is identified as agent-0, and $c_{0}=1$ and $\phi_{0}=0$. Moreover, if the agent-$i$ search cost is the same for all agents $i\in[n]$, then the search cost of the search strategy equals
$\frac{\sqrt{1+k^{2}}}{k}\frac{e^{2\pi k/n}}{\mathcal{G}_{n}}.$
Proof
By appropriately rotating the plane and the search trajectories (without altering them), we may assume that the target is at $(x,0)\in\mathcal{L}_{0}$, where $x>0$. Since $k>0$, there exists $s\in\mathbb{Z}$ such that
$x\in\left(e^{2\pi ks-2\pi k},e^{2\pi ks}\right].$
Now, fix $i\in[n]$. By Lemma 6, agent-$(i+1)$ exposes all points $x\in\left(c_{i}e^{2\pi ks-k\phi_{i}},c_{i}e^{2\pi ks-k\phi_{i+1}}\right],$ where $c_{0}=1$, $\phi_{0}=2\pi$, and agent-$n$ is identified with agent-0. For notational convenience, denote this interval as $\mathcal{I}^{s}_{i}$. By Lemma 5, agent-$(i+1)$ exposes $(x,0)$ at time
$\frac{\ell_{i+1}(2\pi ks-k\phi_{i+1})}{c_{i+1}}=\ell_{0}(2\pi ks-k\phi_{i+1}).$
Thus, the search cost for agent-$(i+1)$ is given by
[…]
concluding the first claim of the lemma.
Now, assume that the agent-$i$ search cost, denoted by $S_{i}$, is invariant for all $i\in[n]$, meaning $S_{i}=S$ for some search cost $S$. Since agent-0 is identified with agent-$n$, and we have $c_{0}=c_{n}=1$ and $\phi_{0}=2\pi$, $\phi_{n}=0$, it follows that
$S^{n}=\prod_{i=1}^{n}S_{i}=\left(\frac{\sqrt{1+k^{2}}}{k}\right)^{n}\frac{1}{\prod_{i=0}^{n-1}c_{i}}\prod_{i=0}^{n-1}\frac{e^{-k\phi_{i+1}}}{e^{-k\phi_{i}}}=\left(\frac{\sqrt{1+k^{2}}}{k}\right)^{n}\frac{1}{\prod_{i=0}^{n-1}c_{i}}e^{2\pi k},$
which establishes the desired formula for $S$. ∎
The next lemma provides a condition on $c\in\mathbb{R}^{n}$ and parameter $k>0$ that ensures $\textsc{MSP}_{n}(c)$ admits a solution using an Off-Set trajectory where all agents expose new points and have the same agent search cost.
Lemma 8
Consider agents $\{c_{i}\}_{i\in[n]}$ satisfying $\min_{i\in[n]}c_{i}\geq e^{-2\pi k/n}\mathcal{G}_{n}$ for some constant $k>0$. For $j\in[n]$, define
$\phi_{j}=\frac{2\pi(n-j)}{n}-\frac{j}{k}\ln\left(\frac{\mathcal{G}_{j}}{\mathcal{G}_{n}}\right).$ (2)
Then, the parameters $\{\phi_{i}\}_{i\in[n]}$ along with $k$ are valid for the Off-Set trajectory $\{g_{i}^{k}\}_{i\in[n]}$, ensuring that all agents expose new points and the agent-$i$ search cost remains independent of $i\in[n]$.
Proof
Fix $k>0$ and angle parameters $\phi_{j}$ as per (2). We use $c_{0}=c_{n}=1$ and $\phi_{n}=0$.
To show that $\{\phi_{i}\}_{i\in[n]}$ along with $k$ are valid parameters for the Off-Set Algorithm 2, we must verify that $\phi_{j}\geq\phi_{j+1}$ for all $j\in[n]$. Equation (2) ensures that $\phi_{0}=0$ and $\phi_{n}=2\pi$. For the remaining parameters, we compute
[…]
where the last inequality follows from $\min_{i\in[n]}c_{i}\geq e^{-2\pi k/n}\mathcal{G}_{n}$.
Next, we prove that all agents expose new points. We must show that
$c_{j}e^{-k\phi_{j}}<c_{j+1}e^{-k\phi_{j+1}}$
for all $j\in[n]$, which, by Lemma 6, is equivalent to all agents exposing new points. Expanding this inequality using (2), we obtain
[…]
where the last inequality follows from $c_{i}\geq e^{-2\pi k/n}\mathcal{G}_{n}$ for all $i\in[n]$, applied at $i=j+1$.
Thus, we have verified that parameters $k>0$ and $\{\phi_{i}\}_{i\in[n]}$ are valid for the Off-Set trajectory $\{g_{i}^{k}\}_{i\in[n]}$ and that all agents expose new points. It remains to show that each agent induces the same search cost.
Let $S_{j}$ denote the agent-$j$ search cost and $S$ the Off-Set trajectory search cost. We show that $S=S_{j}$ if and only if the trajectory parameters satisfy (2). Since we have already established the validity of $\{\phi_{i}\}_{i\in[n]}$ and ensured that all agents expose new points, we apply Lemma 7 to obtain
$S^{j}=\prod_{i=1}^{j}S_{i}=\left(\frac{\sqrt{1+k^{2}}}{k}\right)^{j}\frac{1}{\prod_{i=0}^{j-1}c_{i}}\prod_{i=0}^{j-1}\frac{e^{-k\phi_{i+1}}}{e^{-k\phi_{i}}}=\left(\frac{\sqrt{1+k^{2}}}{k}\right)^{j}\frac{1}{\prod_{i=0}^{j-1}c_{i}}e^{2\pi k-k\phi_{j}}.$
Lemma 7 also provides the formula for $S=S(k,n,c)$. Solving this equation for $\phi_{j}$ yields (2), as required. ∎
4.3 Proof & Discussion of Theorem 4.1
This section is devoted to proving Theorem 4.1, which establishes the correctness of the Off-Set Algorithm 2 when using the parameters selected by Algorithm 3 for input speeds $c\in\mathbb{R}^{n}$. We first prove that Algorithm 3 always returns a valid index $\mu\in[n]$. Then, we verify that the returned parameters are valid inputs to the Off-Set Algorithm 2 and ensure that all agents expose new points. Finally, we compute the search cost for $\textsc{MSP}_{n}(c)$, completing the proof of Theorem 4.1.
Proof(of Theorem 4.1)
We begin by showing that Algorithm 3 always returns some $\mu\in[n]$ when given input $c\in\mathbb{R}^{n}$. Since $l=0$ satisfies the condition in Step 1, the check reduces to verifying $1\geq e^{-2\pi\kappa_{1}}$, which holds because $\kappa_{1}>0$.
Next, we verify that the parameters returned by Algorithm 3 satisfy the conditions required for the Off-Set Algorithm 2. Specifically, let $\mu,\nu,\{\phi_{i}\}_{i\in[\mu]}$ be the returned parameters, where the expansion factor used in Theorem 4.1 is $k=\kappa_{\nu}$. By Step 2, we have $\mu<l$, and by Step 1 and Lemma 8, the parameters $\kappa_{\nu}$ and $\{\phi_{i}\}_{i\in[\mu]}$ are valid for the Off-Set Algorithm 2 using a subset of $\nu$ agents, indexed by $\mu]$. By the same lemma, all agents $i\in[\mu]$ expose new points, and the agent-$i$ search cost is independent of $i$. Thus, applying Lemma 7 and using the definition of $\mathcal{U}_{\nu}$ from Theorem 2.1, we conclude that the induced search cost of the Off-Set algorithm matches the bound given in Theorem 4.1, as required. ∎
To better understand the calculations performed by Algorithm 3, we provide a high-level overview of its operation on input $c\in\mathbb{R}^{n}$. The algorithm effectively selects an index $\mu\in[n]$, meaning that only agents at least as fast as agent-$\mu$ will participate in exposing new points. If agents are sorted in non-increasing order of speed, the participating agents correspond to $\nu=\mu+1$, that is, agents $j$ with $j=0,\ldots,\mu$. The returned angular parameters $\phi_{j}$ are chosen so that the agent search cost is independent of the agent index, ensuring that the points exposed by each agent induce the same cost.
A crucial aspect of Algorithm 3 is that it filters out agents with insufficiently fast speeds. Among all agents with “sufficiently large” speeds, it selects a subset of the strictly fastest agents, denoted as $A$, to minimize the ratio $\mathcal{U}_{|A|}/\mathcal{G}_{|A|}.$
An interesting observation is that Theorem 4.1 generalizes Theorem 2.1, whose proof also serves as a key building block in establishing the former result. In particular, consider $n$ unit-speed agents, meaning $c_{i}=1$ for all $i\in[n]$ is given as input to Algorithm 3. These agents are already sorted in non-increasing order, and their geometric means satisfy $\mathcal{G}_{i}=1$ for every $i\in[n]$. Thus, the largest index $l$ satisfying Step 1 of the algorithm is l=n1l=n-1, as $\mathcal{G}_{n}e^{-2\pi\kappa_{n}/n}=e^{-2\pi\kappa_{n}/n}<1,$ where the inequality holds because $\kappa_{n}>0$. By Proposition 1, the values $\mathcal{U}_{i}$ are strictly decreasing in $i$, so the smallest search cost, as determined by Step 2, is achieved at $\mu=n-1$. Then, in Step 3, the algorithm sets $\nu=\mu+1=n$, and the angular parameters are assigned as $\phi_{i}=\tfrac{2\pi(n-j)}{n},$ for $j\in[n]$, which exactly matches the definition in Algorithm 1.
Thus, Theorem 4.1 not only confirms the validity of the Off-Set algorithm, but also generalizes Theorem 2.1. In particular, the technical proof of Theorem 4.1, which provides upper bound guarantees for arbitrary speed agents, builds directly on the analysis underlying Theorem 2.1 to re-establish the performance guarantee of $\mathcal{U}_{n}$ for the special case of uniform unit-speed agents.
4.4 A Simplified Algorithm for $\mathcal{P}\text{-}\textsc{MSP}_{n}(c)$
This section is devoted to proving Theorem 2.2, that is, we describe a simpler algorithm than $\mathcal{P}\text{-}\textsc{MSP}_{n}(c)$ with the same cost search performance. For this, given input speeds $c\in\mathbb{R}^{n}$ with $c_{0}=1$ (recall we identify agent-0 as agent-$n$, and hence $c_{0}=c_{n}=1$) where the speeds are in non-increasing order, we set
$\ell:=\operatorname*{arg\,min}_{i\in[n]}\mathcal{U}_{i+1}/\mathcal{G}_{i+1}.$ (3)
Recall that $n]=\{0,1,\ldots,n-1\}$. Now consider the fastest $\ell+1$ many agents $\ell+1]$ with speeds $c_{0},c_{1},\ldots,c_{\ell}$. The main technical claim is that index $\ell$, hence speed $c_{\ell}$ of agent-$\ell$, satisfies the condition of Step 1 (to be proven shortly). Hence, we can run the algorithm by setting $l=\ell$. Since $\ell$ was defined as per (3), in Step 2 of Algorithm 3, we must have $\mu=\ell$, and then Theorem 2.2 follows as a corollary of Theorem 4.1.
To summarize, it suffices to show that index $\ell$, as defined in (3), satisfies the condition of Step 1, that is, our goal is to show that
$c_{\ell}\geq e^{-2\pi\kappa_{\ell+1}/\ell}\mathcal{G}_{l}.$ (4)
By the definition (3) of index $\ell$, we have that
$\mathcal{U}_{\ell}/\mathcal{G}_{l}\geq\mathcal{U}_{\ell+1}/\mathcal{G}_{l+1}$
which simplifies, after elementary algebraic manipulations, to the equivalent condition
$c_{\ell}\geq\left(\frac{\mathcal{U}_{\ell+1}}{\mathcal{U}_{\ell}}\right)^{\ell+1}\mathcal{G}_{l}.$
We conclude that a sufficient condition for (4) is that
$\left(\frac{\mathcal{U}_{\ell+1}}{\mathcal{U}_{\ell}}\right)^{\ell}\geq e^{-2\pi\kappa_{\ell+1}/(\ell+1)},$
which we prove next. If $\ell=0$, the inequality holds trivially, because $\kappa_{1}>0$, hence we focus on the case $\ell\geq 1$. To that end we calculate,
$\left(\frac{\mathcal{U}_{\ell+1}}{\mathcal{U}_{\ell}}\right)^{\ell}e^{2\pi\kappa_{\ell+1}/(\ell+1)}=\left(\frac{\sqrt{1+\kappa_{\ell+1}^{2}}/\kappa_{\ell+1}}{\sqrt{1+\kappa_{\ell}^{2}}/\kappa_{\ell}}\right)^{\ell}\frac{e^{2\pi\kappa_{\ell+1}}}{e^{2\pi\kappa_{\ell}}}$
which we would like to show is at least $1$, for all $\ell\geq 1$. This is taken care of by the next technical lemma.
Lemma 9
Let $\kappa_{n}$ be as in Lemma 4. Set
$A(x):=\frac{\sqrt{1+x^{2}}}{x},$
for $x>0$, along with
$H(l):=\left(\frac{A(\kappa_{l+1})}{A(\kappa_{l})}\right)^{l}e^{2\pi\left(\kappa_{l+1}-\kappa_{l}\right)}.$
Then, $H(l)\geq 1$, for all $l\geq 1$.
Proof
We will prove two claims: (a) that $\lim_{l\rightarrow\infty}H(l)=1$, and (b) that $H(l)$ is strictly decreasing in $H(l)$, implying that $H(l)\geq 1$ for all $l\geq 1$.
Regarding (a), we recall Lemma 1 according to which $\kappa_{n}$ is positive, strictly increasing, and strictly concave in $n$. Then, we observe that $A(x)$ is decreasing in $x>0$, as well as that $\lim_{x\rightarrow\infty}A(x)=1$. Moreover, by Lemma 4 it is easy to see that
$\lim_{l\rightarrow\infty}\kappa_{l+1}/\kappa_{l}=1.$
Altogether, this implies also that
$\lim_{l\rightarrow\infty}A(\kappa_{l+1})/A(\kappa_{l})=\lim_{l\rightarrow\infty}\exp\left(2\pi\left(\kappa_{l+1}-\kappa_{l}\right)\right)=1,$
thereby also implying that $\lim_{l\rightarrow\infty}H(l)=1$, as wanted.
Now we treat property (b). We start by setting $T_{l}=\ln\left(H(l)\right),$ so
$T_{l}=l\left[\ln A\left(\kappa_{l+1}\right)-\ln A\left(\kappa_{l}\right)\right]+2\pi\left(\kappa_{l+1}-\kappa_{l}\right).$
We will prove $T_{l}$ is strictly decreasing, and then show $\lim_{l\rightarrow 0}T_{l}=0$, implying this way that $T_{l}\geq 0$, i.e. $H(l)\geq 1$, as wanted.
First we show that $T_{l}$ is strictly decreasing, that is $T_{l+1}<T_{l}$ for all $l\geq 1$. Observe
$T_{l+1}-T_{l}$ $=(l+1)\left(\ln A(\kappa_{l+2})-\ln A(\kappa_{l+1})\right)-l\left(\ln A(\kappa_{l+1})-\ln A(\kappa_{l})\right)+2\pi\left(\kappa_{l+2}-2\kappa_{l+1}+\kappa_{l}\right)$
$=l\cdot X_{1}(l)+X_{2}(l)+2\pi\cdot X_{3}(l),$
where
$X_{1}(l)$ $=\ln A(\kappa_{l+2})-2\ln A(\kappa_{l+1})+\ln A(\kappa_{l}),$
$X_{2}(l)$ $=\ln A(\kappa_{l+2})-\ln A(\kappa_{l+1}),$
$X_{3}(l)$ $=\kappa_{l+2}-2\kappa_{l+1}+\kappa_{l}.$
Next, we argue that $X_{i}(l)<0$, for all $l\geq 1$, for $i=1,2,3$.
We start with $X_{1}(l)$. Set $h_{n}=\ln A\left(\kappa_{n}\right),$ and recall $\kappa_{n}$ is concave and increasing in $n$, while $\ln A(\cdot)$ is convex and decreasing in its argument. Therefore, a standard composition argument shows that $h_{n}$ has negative second finite differences, that is $h_{l+2}-2h_{l+1}+h_{l}<0$ implying that $X_{1}(l)<0$, as wanted.
Then, we treat $X_{2}(l)$. For this, we observe that $A(x)$ is decreasing in $x$, while also $\kappa_{l}$ is increasing in $l$, which implies directly that $A(\kappa_{l+2})/A(\kappa_{l+1})<1$, and hence $X_{2}(l)<0$.
Lastly, $X_{3}(l)<0$ is implied directly by invoking the concavity of $\kappa_{n}$, as argued above. This concludes the proof that $T_{l}$ is strictly decreasing, and hence claim of the lemma. ∎
5 Cone Search with a Unit-Speed Agent
In this section, we design a trajectory for searching for a point within an angle-$\phi$ cone, where $\phi\in(0,\pi)$, using a unit-speed agent. That is, we propose a solution to the problem $\mathcal{C}_{\phi}\text{-}\textsc{MSP}_{1}(1)$ and subsequently prove Theorem 2.3. These results are developed not only for their independent geometric interest, but also to probe the limitations of spiral-based search when agents of different speeds are present. They will later be combined with the spiral-based constructions introduced earlier to obtain improved strategies for the multi-speed point-search problem, providing evidence that spiral-based trajectories, which are optimal for single-agent search, can be suboptimal in the multi-agent, multi-speed setting.
Figure 4: A depiction of the Bouncing Trajectory for $\mathcal{C}_{\phi}\text{-}\textsc{MSP}_{1}(1)$. Angle $\theta$ determines the sequence of points $\Gamma_{i}$, as per Lemma 11, using sequence $q_{i}=\left(\tfrac{\sin\left(\theta\right)}{\sin\left(\theta-\phi\right)}\right)^{i}$.
Algorithm 4 Bouncing Trajectory for $\mathcal{C}_{\phi}\text{-}\textsc{MSP}_{1}(1)$
0: Increasing sequence $\{q_{i}\}_{i\in\mathbb{Z}}$ of reals, with $\lim_{i\rightarrow-\infty}=0$ and $\lim_{i\rightarrow\infty}=\infty$.
 Define points $\Gamma_{2i}=\langle q_{2i},0\rangle$, and $\Gamma_{2i+1}=\langle q_{2i+1},\phi\rangle$, for $i\in\mathbb{Z}$.
Output Trajectory: Piece-wise linear trajectory moving between points $\Gamma_{i}\rightarrow\Gamma_{i+1},~i\in\mathbb{Z}$
Algorithm 4 is illustrated in Figure 4. Next, we derive an upper bound on its search cost.
Lemma 10
The Bouncing Trajectory of Algorithm 4 provides a feasible solution to $\mathcal{C}_{\phi}\text{-}\textsc{MSP}_{1}(1)$. Moreover, the search cost of the algorithm is given by
$\sup_{i\in\mathbb{Z}}\frac{1}{\left\lVert\Gamma_{i}\right\rVert}\sum_{j=-\infty}^{i+2}\left\lVert\Gamma_{j}-\Gamma_{j-1}\right\rVert.$
Proof
First, we observe that the Bouncing Trajectory of Algorithm 4 initially places the agent at the origin, since
$\lim_{i\rightarrow-\infty}\left\lVert\Gamma_{2i}\right\rVert=\lim_{i\rightarrow-\infty}\left\lVert\Gamma_{2i+1}\right\rVert=\lim_{i\rightarrow-\infty}g_{i}=0,$
which implies
$\lim_{i\rightarrow-\infty}\Gamma_{2i}=\lim_{i\rightarrow-\infty}\Gamma_{2i+1}=(0,0).$
Next, consider the sequence of triangles $\{\Delta_{i}\}_{i\in\mathbb{Z}}$, where each $\Delta_{i}$ is formed by the vertices $\Gamma_{i},\Gamma_{i+1},\Gamma_{i+2}$, see also Figure 4. By the properties of the sequence $\{g_{i}\}_{i\in\mathbb{Z}}$, these triangles cover $\mathcal{C}_{\phi}$.
Now, we analyze the worst-case search cost for a target $A\in\Delta_{i}$. The time to expose $A$ is at most the time required for the unit-speed agent to reach $\Gamma_{i+2}$, since by then all points in $\bigcup_{j=-\infty}^{i}\Delta_{j}$ have been exposed. This time is given by
$\sum_{j=-\infty}^{i+2}\left\lVert\Gamma_{j}-\Gamma_{j-1}\right\rVert.$
Furthermore, since $\{g_{j}\}_{j}$ is increasing, we have $\left\lVert A\right\rVert\geq\left\lVert\Gamma_{i}\right\rVert$, confirming that the given expression serves as an upper bound for the search cost. To see that this bound is tight, consider a sequence of targets in the interior of $\Delta_{i}$ converging to $\Gamma_{i}$. ∎
Next, we propose a specific input sequence to Algorithm 4 and derive the induced search cost for $\mathcal{C}_{\phi}\text{-}\textsc{MSP}_{1}(1)$.
Lemma 11
For any $\phi\in(0,\pi)$ and $\theta\in(\phi,\pi/2+\phi/2)$, let
$\alpha$ $:=\tfrac{\sin\left(\phi\right)}{\sin\left(\theta\right)},$
$\beta$ $:=\tfrac{\sin\left(\theta\right)}{\sin\left(\theta-\phi\right)}.$
Then, the sequence $q_{i}=\beta^{i},i\in\mathbb{Z},$ is a valid input to Algorithm 4, inducing a search cost of
$\tfrac{\alpha\beta^{3}}{\beta-1}$
for $\mathcal{C}_{\phi}\text{-}\textsc{MSP}_{1}(1)$.
Proof
Fix $\phi\in(0,\pi)$. We first show that $\beta>1$, ensuring that the sequence is well-defined. Using the trigonometric identity
$\sin\left(\theta\right)-\sin\left(\theta-\phi\right)=2\cos\left(\theta-\phi/2\right)\sin\left(\phi/2\right),$
and noting that $\sin\left(\phi/2\right)>0$ for $\phi\in(0,\pi)$, we analyze the term $\cos\left(\theta-\phi/2\right)$. Since $\theta\in(\phi,\pi/2+\phi/2)$, it follows that
$\theta-\phi/2\in(\phi/2,\pi/2)\subseteq(0,\pi/2),$
which implies $\cos\left(\theta-\phi/2\right)>0$. Also, since $\theta-\phi\in(0,\pi/2-\phi/2)$, we conclude that $\sin\left(\theta-\phi\right)>0$, establishing $\beta=\sin\left(\theta\right)/\sin\left(\theta-\phi\right)>1$.
Next, we compute the points $\Gamma_{i}$ in Algorithm 4 and derive the search cost using Lemma 10. The parameter $\beta=\beta(\phi,\theta)$ defines the sequence $\{g_{j}\}_{j}$, which uniquely determines the points $\{\Gamma_{j}\}_{j}$ and vice versa. We now provide a geometric and inductive definition of these points. For reference, Figure 4 illustrates the construction and highlights the role of the angle $\theta$.
Setting $g_{0}=1$, we define $\Gamma_{0}=(1,0)\in\mathcal{L}_{0}$. For each $i\in\mathbb{Z}$, we require that $\Gamma_{i}$ and $\Gamma_{i+1}$ lie on different extreme rays of $\mathcal{C}_{\phi}$ and that the vertex angle at $\Gamma_{i}$ in the triangle $O\Gamma_{i}\Gamma_{i+1}$ (where $O$ is the origin) is $\pi-\theta$. This definition uniquely determines $\Gamma_{i}$ and thus $g_{i}$ for all $i\in\mathbb{Z}$.
Now, we compute $g_{i}$ and $d_{i}:=\left\lVert\Gamma_{i}-\Gamma_{i-1}\right\rVert$. Applying the Law of Sines to triangle $O\Gamma_{i}\Gamma_{i+1}$, and using $\left\lVert\Gamma_{i}\right\rVert=g_{i}$, we obtain
$\frac{d_{i+1}}{\sin\left(\phi\right)}=\frac{g_{i+1}}{\sin\left(\pi-\theta\right)}=\frac{g_{i+1}}{\sin\left(\theta\right)},$
since $\theta\in(0,\pi)$. Therefore,
$d_{i+1}=\tfrac{\sin\left(\phi\right)}{\sin\left(\theta\right)}g_{i+1}=\alpha g_{i+1}.$
Similarly, using the fact that the angle at $O$ is $\phi$, we find that the angle at $\Gamma_{i+1}$ is $\theta-\phi\in(0,\pi/2)$. Applying the Law of Sines again,
$\frac{g_{i+1}}{\sin\left(\pi-\theta\right)}=\frac{g_{i+1}}{\sin\left(\theta-\phi\right)},$
which implies
$g_{i+1}=\tfrac{\sin\left(\theta\right)}{\sin\left(\theta-\phi\right)}g_{i}=\beta g_{i}.$
Since $g_{0}=1$, it follows that $g_{i}=\beta^{i}$, and consequently, $d_{i}=\alpha\beta^{i}$.
Using Lemma 10, the search cost of Algorithm 4 is
$\sup_{i\in\mathbb{Z}}\frac{1}{\left\lVert\Gamma_{i}\right\rVert}\sum_{j=-\infty}^{i+2}\left\lVert\Gamma_{j}-\Gamma_{j-1}\right\rVert=\sup_{i\in\mathbb{Z}}\frac{1}{g_{i}}\sum_{j=-\infty}^{i+2}d_{j}=\sup_{i\in\mathbb{Z}}\frac{\alpha}{\beta^{i}}\sum_{j=-\infty}^{i+2}\beta^{j}=\sup_{i\in\mathbb{Z}}\frac{\alpha}{\beta^{i}}\frac{\beta^{i+3}}{\beta-1}=\frac{\alpha\beta^{3}}{\beta-1}.$
Thus, the search cost is independent of $i$, completing the proof. ∎
We now summarize our findings with the proof of Theorem 2.3.
Proof(of Theorem 2.3)
Given $\phi\in(0,\pi)$, we apply Algorithm 4 to $\mathcal{C}_{\phi}\text{-}\textsc{MSP}_{1}(1)$ with input sequence
$q_{i}=\left(\frac{\sin\left(\theta\right)}{\sin\left(\theta-\phi\right)}\right)^{i},i\in\mathbb{Z}.$
By Lemma 11, if $\theta\in(\phi,\pi/2+\phi/2)$, the sequence is valid for the algorithm, and the resulting search cost is
$\frac{\sin\left(\phi\right)}{\sin\left(\theta\right)}\cdot\frac{\beta^{3}}{\beta-1}=\frac{\sin\left(\phi\right)}{2\sin\left(\frac{\phi}{2}\right)}\cdot\left(\frac{\sin\left(\theta\right)}{\sin\left(\theta-\phi\right)}\right)^{2}\cdot\frac{1}{\cos\left(\theta-\phi/2\right)}.$
This represents the search cost of the algorithm.
Now, for every $\theta\in(\phi,\pi/2+\phi/2)$, the search cost function is positive and continuous. Moreover, it tends to $\infty$ as $\theta$ approaches either $\phi^{+}$ or $(\pi/2+\phi/2)^{-}$. By the Extreme Value Theorem, the function attains a minimum within the open interval $(\phi,\pi/2+\phi/2)$. ∎
6 Conic Complement Search with a Unit-Speed Agent
In this section, we design a trajectory for searching a point in the angle-$\phi$ conic complement, where $\phi\in(0,\pi)$, using a unit-speed agent. That is, we propose a solution to the problem $\mathcal{W}_{\phi}\text{-}\textsc{MSP}_{1}(1)$, and subsequently prove Theorem 2.4.
To describe the trajectory for searching $\mathcal{W}_{\phi}$, we first develop some intuition for why this can be done more efficiently than searching the entire plane $\mathcal{P}$ with a unit-speed agent, even when $\phi\in(0,\pi)$ is small. A key aspect in analyzing Algorithm 1 for $\mathcal{P}\text{-}\textsc{MSP}_{1}(1)$ is that the search cost is independent of the target’s angular coordinate. For this reason, and without loss of generality, we fix a reference ray $\mathcal{L}_{0}$ and perform worst-case analysis under the assumption that the target lies on $\mathcal{L}_{0}$.
Let $f_{0}^{k}(t)$ denote the trajectory of Algorithm 1, applicable to $\mathcal{P}\text{-}\textsc{MSP}_{1}(1)$. This trajectory intersects $\mathcal{L}_{0}$ at the points $X_{s}=f_{0}^{k}(2\pi s)$ and intersects the ray $\mathcal{L}_{\phi}$ at the points $Z_{s}=f_{0}^{k}(2\pi s+\phi)$, where $s\in\mathbb{Z}$, see also Figure 5. To adapt $f_{0}^{k}(t)$ for searching $\mathcal{W}_{\phi}$ (i.e., the plane excluding the cone $\mathcal{C}_{\phi}$), we can modify the trajectory by replacing the curved segment between $X_{s}$ and $Z_{s}$ with the shorter straight line segment $X_{s}Z_{s}$, while preserving the original trajectory between $Z_{s}$ and $X_{s+1}$. With this modification, every point in $\mathcal{W}_{\phi}$ is reached strictly faster. However, the search cost becomes dependent on the target’s location. Since the search cost is defined as the supremum over all possible target positions in the unbounded domain $\mathcal{W}_{\phi}$, this modification does not necessarily reduce the worst-case cost.
In the following, we explore this idea in more detail by properly defining the sequences of points $X_{s}$ and $Z_{s}$, ensuring that the search cost remains independent of the target placement. This, in turn, simplifies the analysis and allows us to demonstrate a strict improvement over searching the full plane.
Figure 5: A depiction of the Shortcut Spiral Trajectory to $\mathcal{W}_{\phi}\text{-}\textsc{MSP}_{1}(1)$. The shaded cone corresponds to $\mathcal{C}_{\phi}$ with extreme rays $\mathcal{L}_{0},\mathcal{L}_{\phi}$, while the shortcut is represented by the dashed arrow within the cone.
Algorithm 5 Shortcut Spiral Trajectory to $\mathcal{W}_{\phi}\text{-}\textsc{MSP}_{1}(1)$
0:$k>0$, $1<\lambda<e^{2\pi k}$.
 For $s\in\mathbb{Z}$, set $X_{s}:=\langle e^{2\pi ks},0\rangle$, and $Y_{s}:=\langle\lambda e^{2\pi ks},\phi\rangle$.
 Set $\ell:=\ln(\lambda)$, $\alpha:=\tfrac{2\pi k-\ell}{2\pi-\phi}$, and $\beta_{s}:=2\pi ks+\ell-\alpha\phi$.
 Set $h_{s}(t):=\alpha t+\beta_{s}$.
Output Trajectory: Line segment $X_{s}\rightarrow Y_{s}$, followed by $p_{s}(t)=\langle e^{h_{s}(t)},t\rangle$ for $t\in[{\phi,2\pi}]$, and $s\in\mathbb{Z}$
As before, the conditions $k>0$ and $1<\lambda<e^{2\pi k}$ ensure that the trajectory exposes all points in $\mathcal{W}_{\phi}$. In particular, the inequality $\lambda<e^{2\pi k}$ implies that $\alpha>0$. Choosing $\lambda=e^{k\phi}$ recovers the original description of the cross-cut algorithm, where the spiral follows the trajectory of Algorithm 1 while remaining within $\mathcal{W}_{\phi}$.
Lemma 12
For every $\phi\in(0,\pi/2)$, Algorithm 5 provides a feasible solution to $\mathcal{W}_{\phi}\text{-}\textsc{MSP}_{1}(1)$.
Proof
We begin by showing that the trajectory is continuous. The trajectory is defined inductively over $s\in\mathbb{Z}$. For each fixed $s$, it moves from $X_{s}\in\mathcal{L}_{0}$ to $Y_{s}\in\mathcal{L}_{\phi}$, followed by a spiral segment $p_{s}(t)$ for $t\in[\phi,2\pi]$, see also Figure 5. Observe that
$h_{s}(\phi)$ $=\alpha\phi+\beta_{s}=\alpha\phi+2\pi ks+\ell-\alpha\phi=2\pi ks+\ell,$
$h_{s}(2\pi)$ $=\alpha 2\pi+\beta_{s}=\alpha(2\pi-\phi)+2\pi ks+\ell=2\pi k(s+1).$
This implies
$\left\lVert p_{s}(\phi)\right\rVert=e^{h_{s}(\phi)}=\lambda e^{2\pi ks}=\left\lVert Y_{s}\right\rVert,\quad\left\lVert p_{s}(2\pi)\right\rVert=e^{h_{s}(2\pi)}=e^{2\pi k(s+1)}=\left\lVert X_{s+1}\right\rVert.$
Hence, $p_{s}(\phi)=Y_{s}$ and $p_{s}(2\pi)=X_{s+1}$, proving that the trajectory is continuous.
Next, we verify feasibility. Since $k>0$, we have $\lim_{s\to-\infty}X_{s}=(0,0)$, so the agent starts at the origin. It remains to show that every point in $\mathcal{W}_{\phi}$ is eventually exposed.
Let $A=\langle x,\psi\rangle$ with $\psi\in(\phi,2\pi)$ and $x>0$. For each $s$,
$h_{s}(\psi)=\alpha\psi+\beta_{s}=\alpha\psi+2\pi ks+\ell-\alpha\phi.$
Since $k>0$, $h_{s}(\psi)$ becomes arbitrarily large as $s\to\infty$. Hence, for sufficiently large $s$, we have $x<e^{h_{s}(\psi)}$, implying $p_{s}(\psi)$ exposes $A$. ∎
It follows that for each $s\in\mathbb{Z}$, the set of targets exposed by $p_{s}(t)$, for $t\in(\phi,2\pi)$, forms a nested family in $s$, and their union covers $\mathcal{W}_{\phi}$. Define the wedge $W_{s}\subseteq\mathcal{W}_{\phi}$ as the set of targets first exposed by $p_{s+1}(t)$ for some $t\in(\phi,2\pi)$. The wedges $\{W_{s}\}_{s}$ are disjoint, and by Lemma 12, we have $\bigcup_{s}W_{s}=\mathcal{W}_{\phi}$.
We now analyze the search cost of Algorithm 5. The following observation will be useful.
Lemma 13
For every $s\in\mathbb{Z}$ and $t_{1},t_{2}\in\mathbb{R}$, we have
$\int_{t_{1}}^{t_{2}}\left\lVert p_{s}^{\prime}(t)\right\rVert\,\mathrm{d}t=\frac{\sqrt{1+\alpha^{2}}}{\alpha}e^{2\pi ks}\left(e^{h_{0}(t_{2})}-e^{h_{0}(t_{1})}\right),$
and
$\left\lVert X_{s}Y_{s}\right\rVert=e^{2\pi ks}\left\lVert X_{0}Y_{0}\right\rVert,$
where $\alpha$, $p_{s}(t)$, $X_{s}$, and $Y_{s}$ are as defined in Algorithm 5.
Proof
By Lemma 2, and since $\alpha>0$,
$\int_{t_{1}}^{t_{2}}\left\lVert p_{s}^{\prime}(t)\right\rVert\,\mathrm{d}t=\frac{\sqrt{1+\alpha^{2}}}{\alpha}\left(e^{h_{s}(t_{2})}-e^{h_{s}(t_{1})}\right)=\frac{\sqrt{1+\alpha^{2}}}{\alpha}e^{2\pi ks}\left(e^{h_{0}(t_{2})}-e^{h_{0}(t_{1})}\right),$
using the identity $h_{s}(t)=2\pi ks+h_{0}(t)$. Also, in Cartesian coordinates, $X_{s}=e^{2\pi ks}(1,0)$ and
$Y_{s}=\lambda e^{2\pi ks}(\cos\left(\phi\right),\sin\left(\phi\right)),$
so
$\left\lVert X_{s}Y_{s}\right\rVert=e^{2\pi ks}\left\lVert(1,0)-\lambda(\cos\left(\phi\right),\sin\left(\phi\right))\right\rVert=e^{2\pi ks}\left\lVert X_{0}Y_{0}\right\rVert.$
We now compute the time required to expose a point on $\mathcal{L}_{\psi}$, where $\phi\in[\phi,2\pi]$, which guides the optimal choice of $k$ and $\lambda$ to minimize the search cost.
Lemma 14
For every $\psi\in[\phi,2\pi]$ and $s\in\mathbb{Z}$, any target in $W_{s}\cap\mathcal{L}_{\psi}$ is exposed by $p_{s+1}(\psi)$ at time
$e^{2\pi k(s+1)}\left(\frac{\sqrt{1+\alpha^{2}}}{\alpha}\lambda e^{\alpha(\psi-\phi)}+\frac{e^{2\pi k}}{e^{2\pi k}-1}\left(\left\lVert X_{0}Y_{0}\right\rVert-\frac{\sqrt{1+\alpha^{2}}}{\alpha}(\lambda-1)\right)\right).$
Proof
We have $p_{s+1}(\psi)\in\mathcal{L}_{\psi}$ with $\left\lVert p_{s+1}(\psi)\right\rVert=e^{h_{s+1}(\psi)}=e^{2\pi k}e^{h_{s}(\psi)},$ and by definition of $W_{s}$, this is at least $\left\lVert A\right\rVert$ for all $A\in W_{s}\cap\mathcal{L}_{\psi}$. By Lemma 13, the time to reach $p_{s+1}(\psi)$ is
$\sum_{i=-\infty}^{s+1}\left\lVert X_{i}Y_{i}\right\rVert+\sum_{i=-\infty}^{s}\int_{\phi}^{2\pi}\left\lVert p_{i}^{\prime}(t)\right\rVert\,\mathrm{d}t+\int_{\phi}^{\psi}\left\lVert p_{s+1}^{\prime}(t)\right\rVert\,\mathrm{d}t$
$=\left\lVert X_{0}Y_{0}\right\rVert\sum_{i=-\infty}^{s+1}e^{2\pi ki}+\frac{\sqrt{1+\alpha^{2}}}{\alpha}\left(\left(e^{h_{0}(2\pi)}-e^{h_{0}(\phi)}\right)\sum_{i=-\infty}^{s}e^{2\pi ki}+e^{2\pi k(s+1)}\left(e^{h_{0}(\psi)}-e^{h_{0}(\phi)}\right)\right).$
Using $h_{0}(t)=\alpha(t-\phi)+\ell$, $h_{0}(\phi)=\ell$, and $h_{0}(2\pi)=2\pi k$, and simplifying gives the claimed expression. ∎
We now show that we can choose $k$ and $\lambda$ to make the search cost independent of the target’s location.
Lemma 15
For $\phi\in(0,\pi)$ and any $\lambda>1$, define
$k_{\lambda}:=\frac{\ln\lambda}{2\pi}+\frac{\lambda-1}{\sqrt{2\lambda(1-\cos\left(\phi\right))}}\left(1-\frac{\phi}{2\pi}\right),$
and
$\alpha_{\lambda}:=\frac{2\pi k_{\lambda}-\ln\lambda}{2\pi-\phi}.$
Then, $(k_{\lambda},\lambda)$ is a valid input to Algorithm 5 for $\mathcal{W}_{\phi}\text{-}\textsc{MSP}_{1}(1)$, and the search cost equals
$\frac{\sqrt{1+\alpha_{\lambda}^{2}}}{\alpha_{\lambda}}e^{2\pi k_{\lambda}}.$
Proof
Since $\lambda>1$, it follows that $k_{\lambda}>\frac{\ln\lambda}{2\pi}>0$, so $2\pi k_{\lambda}>\ln\lambda$, and hence $\lambda<e^{2\pi k_{\lambda}}$, satisfying the algorithm’s conditions.
Let $\psi\in(\phi,2\pi)$ and $A\in W_{s}\cap\mathcal{L}_{\psi}$. By Lemma 14, $A$ is exposed by $p_{s+1}(\psi)$, and since
$\left\lVert A\right\rVert\geq\left\lVert p_{s}(\psi)\right\rVert=\lambda e^{2\pi ks}e^{\alpha(\psi-\phi)},$
the search cost is at most
$\frac{\sqrt{1+\alpha^{2}}}{\alpha}e^{2\pi k}+\frac{e^{4\pi k}}{\lambda(e^{2\pi k}-1)e^{\alpha(\psi-\phi)}}\left(\left\lVert X_{0}Y_{0}\right\rVert-\frac{\sqrt{1+\alpha^{2}}}{\alpha}(\lambda-1)\right).$
Elementary calculations show that $\left\lVert X_{0}Y_{0}\right\rVert=1+\lambda^{2}-2\lambda\cos\left(\phi\right)$. Recalling that $\alpha=\tfrac{2\pi k-\ell}{2\pi-\phi}$, and by applying simple algebraic manipulations, we see that the choice of $k=k_{\lambda}$ is the one that makes $\left\lVert X_{0}Y_{0}\right\rVert-\tfrac{\sqrt{1+\alpha^{2}}}{\alpha}(\lambda-1)$ equal to 0, that is, it makes the search cost independent of the placement of the target. ∎
We are now ready to prove Theorem 2.4, which follows from Lemma 15 by showing that the search cost admits a minimum for every $\phi\in(0,\pi)$ when the parameter $\lambda$ is restricted to values strictly greater than $1$.
Proof(of Theorem 2.4)
By Lemma 15, for all $\lambda>1$, and setting $k=k_{\lambda}$, the search cost of Algorithm 5 is
$p(\lambda):=\frac{\sqrt{1+\alpha_{\lambda}^{2}}}{\alpha_{\lambda}}e^{2\pi k_{\lambda}},$
where $\alpha_{\lambda}=\frac{2\pi k_{\lambda}-\ln\lambda}{2\pi-\phi}$. We now simplify $p(\lambda)$ and show that for every $\phi\in(0,\pi)$, this function attains a minimum over $\lambda>1$.
First, observe that $p(\lambda)$ is continuous for all $\lambda>1$, and since $\alpha_{\lambda}>0$, we have $p(\lambda)\geq 0$. Substituting the expression for $\alpha_{\lambda}$ from Lemma 15,
$\alpha_{\lambda}=\frac{\lambda-1}{\sqrt{2\lambda(1-\cos\left(\phi\right))}},$
we obtain the simplified form
$p(\lambda)=\frac{\sqrt{\lambda^{2}-2\lambda\cos(\phi)+1}}{\lambda-1}e^{2\pi k_{\lambda}}.$
We now examine the behavior at the boundaries. As $\lambda\to\infty$, we have $\lim_{\lambda\to\infty}k_{\lambda}=\infty$, and
$\lim_{\lambda\to\infty}\frac{\sqrt{\lambda^{2}-2\lambda\cos(\phi)+1}}{\lambda-1}=1,$
so $\lim_{\lambda\to\infty}p(\lambda)=\infty$. As $\lambda\to 1^{+}$, we have $\lim_{\lambda\to 1^{+}}k_{\lambda}=0$, but
$\lim_{\lambda\to 1^{+}}\frac{\sqrt{\lambda^{2}-2\lambda\cos(\phi)+1}}{\lambda-1}=\infty,$
implying $\lim_{\lambda\to 1^{+}}p(\lambda)=\infty$. Thus, the function $p(\lambda)$ is continuous and non-negative on $(1,\infty)$ and diverges at both endpoints. By the Extreme Value Theorem, $p(\lambda)$ attains a minimum in the open interval $(1,\infty)$. ∎
Theorem 2.4 will be used in the next section to establish the performance of Algorithm 2 for $\mathcal{P}\textrm{-}\textsc{MSP}_{2}(1,c)$. Algorithm 5 will be utilized by the speed-$1$ agent. Recall that for each $\phi$, Algorithm 5 is parameterized by $\lambda,k$, whereas the optimal parameters are set to $\lambda=\lambda(\phi)$ and $k_{\lambda}=k_{\lambda}(\phi)$ by Theorem 2.4. At the same time, Algorithm 5 emerged by modifying Algorithm 1 which corresponds to choosing $\lambda=e^{k\phi}$. We are therefore motivated to explore for each $\phi$ how $e^{k_{\lambda}\phi}$ compares to $\lambda$. This is done in Figure 6.
(a) Comparison between $e^{k_{\lambda}\phi}$ and $\lambda$, for the optimal values of $\lambda(\phi),k_{\lambda}(\phi)$ as functions of $\phi$. The plot shows $e^{k_{\lambda}\phi}-\lambda$, as a function of $\phi$.
(b) The plot of the optimal $k_{\lambda}=k_{\lambda}(\phi)$ as it is used by Algorithm 5.
Figure 6: Plot of parameters $\lambda=\lambda(\phi)$ and $k_{\lambda}=k_{\lambda}(\phi)$ as computed by Theorem 2.4 and used in Algorithm 5 for searching $\mathcal{W}_{\phi}$, $\phi\in(0,\pi)$. The horizontal axis in both plots corresponds to the angular parameter $\phi$.
7 Sub-Optimality of the Spiral-Type Off-Set Algorithm
In this section, we demonstrate that the best performance achieved by the logarithmic-spiral-type Off-Set Algorithm 2 for planar search, as described in Theorem 2.2, can be strictly improved when the speed of one of the agents is sufficiently small. We illustrate this by focusing on the special case of two agents searching the plane, that is, the problem $\mathcal{P}\text{-}\textsc{MSP}_{2}(1,c)$, where $c\in(0,1)$. This establishes Theorems 2.5 and 2.6. We begin with an observation.
Proposition 2
The optimal search cost for $\mathcal{P}\text{-}\textsc{MSP}_{2}(1,c)$ with $c<1/\mathcal{U}_{1}\approx 0.0578391$ is $\mathcal{U}_{1}\approx 17.28935$.
Proof
The optimal search cost for $\textsc{MSP}_{2}(1,c)$ is at most $\mathcal{U}_{1}$, since the unit-speed agent can simulate the optimal solution to $\textsc{MSP}_{1}(1)$. We now show that, when $c<1/\mathcal{U}_{1}$, the speed-$c$ agent cannot be the first to expose any point in an optimal solution.
Suppose, for contradiction, that an optimal algorithm for $\textsc{MSP}_{2}(1,c)$ has the speed-$c$ agent first exposing a point $A$. Consider a unit-speed agent following the same trajectory. The unit-speed agent must take at least 1 unit of time to expose $A$, so the speed-$c$ agent, moving $1/c$ times slower, takes at least $1/c$ time to reach $A$. Hence, the search cost for exposing $A$ would be at least $1/c>\mathcal{U}_{1}$, contradicting the assumption that the algorithm is optimal. ∎
Proposition 2 is a negative result, as it shows that slow agents with speed $c<1/\mathcal{U}_{1}$ cannot help reduce the optimal search cost below $\mathcal{U}_{1}$ in the problem $\mathcal{P}\text{-}\textsc{MSP}_{2}(1,c)$. However, can the Off-Set logarithmic-spiral Algorithm 2 make use of such slow-moving agents? The following immediate corollary of Theorem 2.2 addresses this.
Corollary 2
The performance of the Off-Set Algorithm 2 for $\mathcal{P}\text{-}\textsc{MSP}_{2}(1,c)$ equals $\mathcal{U}_{1}$ if $c\leq(\mathcal{U}_{2}/\mathcal{U}_{1})^{2}\approx 0.266687$, and $\mathcal{U}_{2}/\sqrt{c}$ if $c>(\mathcal{U}_{2}/\mathcal{U}_{1})^{2}$.
Next, we formally demonstrate how to strictly improve the performance of the Off-Set Algorithm 2 for $\mathcal{P}\text{-}\textsc{MSP}_{2}(1,c)$ when
$1/\mathcal{U}_{1}<c<\left(\mathcal{U}_{2}/\mathcal{U}_{1}\right)^{2}.$
Moreover, this improvement extends to speeds slightly exceeding the threshold $\left(\mathcal{U}_{2}/\mathcal{U}_{1}\right)^{2}$, which marks the limit beyond which Algorithm 2 ceases to be optimal for $\mathcal{P}\text{-}\textsc{MSP}_{2}(1,c)$. In this extended range, the speed-$c$ agent exposes new points while incurring a strictly lower search cost than $\mathcal{U}_{1}$, though further improvements are still possible.
The Off-Set Algorithm 2 is spiral based in the sense that both agents are tasked with covering all directions, and the analysis enforces a shared worst case exposure across agents. In the regime of Corollary 2, this forces a dichotomy: either the slow agent does not help at all, or it participates but drives the bound to the spiral expression $\mathcal{U}_{2}/\sqrt{c}$. The hybrid strategy below avoids this global coupling by partitioning the plane into two geometrically different regions, and assigning each agent only the region it can search efficiently. The speed-$c$ agent is restricted to a cone, where directional search is effective, while the unit-speed agent covers the conic complement using the wedge search bound, and this separation is what enables an improvement over two spirals for an intermediate range of speeds.
We begin by describing an alternative algorithm for $\mathcal{P}\text{-}\textsc{MSP}_{2}(1,c)$.
Algorithm 6 Cone-Wedge Hybrid Algorithm for $\mathcal{P}\text{-}\textsc{MSP}_{2}(1,c)$
[…]
 Sequence $\{q_{i}\}_{i\in\mathbb{Z}}$ defined as in Lemma 11, and fixed to $q_{i}=q_{i}(\phi)$ by Theorem 2.3.
 Parameters $\lambda,k_{\lambda}$ defined in Lemma 15, and fixed to $\lambda=\lambda(\phi)$, $k_{\lambda}=k_{\lambda}(\phi)$ by Theorem 2.4.
Output Trajectories: The speed-$c$ agent searches $\mathcal{C}_{\phi}$ using Algorithm 4 with input sequence $\{q_{i}\}_{i\in\mathbb{Z}}$.The unit-speed agent searches $\mathcal{W}_{\phi}$ using Algorithm 5 with parameters $\lambda$ and $k_{\lambda}$.
We now demonstrate the effectiveness of Algorithm 6 by proving Theorem 2.5.
Proof(of Theorem 2.5)
We analyze the performance of Algorithm 6 for input $\phi\in(0,\pi)$. The plane is partitioned into the cone $\mathcal{C}_{\phi}$, searched by the speed-$\gamma(\phi)$ agent, and the wedge $\mathcal{W}_{\phi}$, searched by the unit-speed agent. By Theorem 2.3, the search cost for points in $\mathcal{C}_{\phi}$ is
$\cos\left(\phi/2\right)\cdot\frac{\mathcal{F}_{\phi}}{\gamma(\phi)},$
i.e., the unit-speed cost scaled by $1/\gamma(\phi)$. Similarly, Theorem 2.4 gives a search cost of $\mathcal{R}_{\phi}$ for points in $\mathcal{W}_{\phi}$. By the definition of $\gamma(\phi)$, these two expressions are equal, so the overall search cost is $\mathcal{R}_{\phi}$, as claimed. ∎
Theorem 2.6 now follows as a numerical corollary of Theorem 2.5. For it’s proof, we numerically compute the pairs $(\gamma(\phi),\mathcal{R}_{\phi})$ for $\phi\in(0,\pi/2)$, where a speed-$\gamma(\phi)$ agent searches the cone $\mathcal{C}_{\phi}$, incurring a corresponding search cost $\mathcal{R}_{\phi}$. We compare these values to the upper bound $\mathcal{U}_{2}/\sqrt{c}$ given in Theorem 2.2, as shown in Figure 1. Since our goal is to optimize performance for small values of $c$, the approach is valid as long as the range of $\gamma(\phi)$ contains the interval
$\left(1/\mathcal{U}_{1},\left(\mathcal{U}_{2}/\mathcal{U}_{1}\right)^{2}\right)\approx(0.0578391,0.266687).$
Numerical calculations which are depicted in Figure 7 confirm that this interval is indeed fully covered.
Figure 7: Plot of $\gamma(\phi)=\cos\left(\phi/2\right)\tfrac{\mathcal{F}_{\phi}}{\mathcal{R}_{\phi}}$ against $\phi$, where $\mathcal{F}_{\phi}$ and $\mathcal{R}_{\phi}$ are defined in Theorems 2.3 and 2.4, respectively.
8 Discussion
We studied point search in the plane using multi-speed agents, focusing on extensions of spiral-based strategies that are known to be optimal in the single-agent, unit-speed setting. While spirals provide a natural baseline, our results show that they need not be optimal once agents have heterogeneous speeds. In particular, we identify a concrete regime in which spiral-based offset strategies become suboptimal, providing the first rigorous evidence of a qualitative separation between single-speed and multi-speed search.
The quantitative improvement obtained by the conic-complement hybrid strategy is modest and confined to a narrow range of speed ratios. We do not claim that this hybrid is close to optimal, and we currently lack intuition about the structure of globally optimal strategies in this regime. Rather, the hybrid construction serves to demonstrate that assigning geometrically distinct tasks to agents of different speeds can outperform frameworks in which all agents are required to cover all directions.
At the level of subproblems, stronger statements may be possible. For directional (cone) search, the cone-search strategy appears to capture the essential geometric difficulty of the problem, and we conjecture that it is optimal in that setting. For $n$ agents of equal unit speed, we likewise conjecture that spiral-based strategies remain optimal, although establishing this lies outside the scope of the present work. Finally, for sufficiently large speed ratios, we conjecture that the offset spiral strategy analyzed here becomes optimal again, reflecting the diminishing influence of slower agents.
Randomized strategies define a largely orthogonal direction. Randomizing the global orientation of a deterministic trajectory removes the adversary’s ability to select a worst-case target angle, so that the expected performance coincides with the average-case performance of the corresponding deterministic strategy under a uniformly random target direction. A systematic comparison between deterministic and randomized strategies in the multi-speed planar search model remains open.
Several further questions arise naturally. Can matching lower bounds be established for restricted classes of speed distributions? Does a similar separation between spiral-based and non-spiral strategies arise in shoreline search with heterogeneous agents? How should strategies adapt when agent speeds are unknown or time-varying, or when search is conducted in bounded domains or environments with obstacles?